### v1

# A Note on Output Length of One-Way State Generators

Abstract
We study output length of one-way state generators (OWSGs) and their weaker variants.
  • Standard OWSGs.
    Recently, Cavalar et al. (arXiv:2312.08363) give OWSGs with $m$-qubit outputs for any $m=\omega(\log\lambda)$, where $\lambda$ is the security parameter, and conjecture that there do not exist OWSGs with $O(\log\log\lambda)$-qubit outputs. We prove their conjecture in a stronger manner by showing that there do not exist OWSGs with $O(\log\lambda)$-qubit outputs. This means that their construction is optimal in terms of output length.
  • Constant-advantage OWSGs.
    Let $\epsilon$-OWSGs be a parameterized variant of OWSGs where a quantum polynomial-time adversary’s advantage is at most $\epsilon$. For any constant $\epsilon>0$, we construct $\epsilon$-OWSGs with $O(\log\log\lambda)$-qubit outputs assuming the existence of subexponentially secure OWFs. We show that this is almost tight by proving that there do not exist $O(1)$-OWSGs with $(\log\log\lambda)/2+O(1)$-qubit outputs.
  • Weak OWSGs.
    We refer to $(1-1/{\mathrm{poly}}(\lambda))$-OWSGs as weak OWSGs. We construct weak OWSGs with $m$-qubit outputs for any $m=\omega(1)$ assuming the existence of exponentially secure injective OWFs with linear expansion. We show that this is tight by proving that there do not exist weak OWSGs with $O(1)$-qubit outputs.
1 Introduction
In classical cryptography, the existence of one-way functions (OWFs) is the minimum assumption, because many primitives (such as secret-key encryption (SKE), commitments, zero-knowledge, pseudorandom generators (PRGs), pseudorandom functions (PRFs), and digital signatures) are equivalent to OWFs, and almost all cryptographic primitives imply OWFs [LR86, IL89, ILL89]. On the other hand, in quantum cryptography where quantum computing and quantum communications are available, recent studies suggest that OWFs would not be the minimum assumption, and many new replacements of OWFs have been introduced. For example, pseudorandom state generators (PRSGs) [JLS18], one-way state generators (OWSGs) [MY22b, MY22a], and EFI [BCQ23] are quantum analogue of PRGs, OWFs, and EFID [Gol90], respectively. [^1] EFI is a pair of efficiently generatable states that are statistically distinguishable but computationally not. EFID is its classical version: a pair of efficiently samplable classical distributions that are statistically distinguishable but computationally not. They seem to be weaker than OWFs [Kre21, KQST23, LMW23], yet can be used to construct several useful primitives such as private-key quantum money, SKE, commitments, multiparty computations, and digital signatures [MY22b, AQY22, BCQ23, Yan22].
In this paper, we focus on OWSGs, which are a quantum analog of OWFs.222 There are a pure state output version [MY22b] and mixed state output version [MY22a] of OWSGs. We consider OWSGs with mixed state outputs as a default definition, but all of our positive results actually have pure state outputs. We adopt the mixed state version to make our negative results stronger. A OWSG is a quantum polynomial-time (QPT) algorithm that takes a bit string $k$ as input and outputs a quantum state $\phi_{k}$. The security is, roughly speaking, that no QPT adversary can find “preimage” of $\phi_{k}$. There are two differences from classical OWFs. First, because quantum states cannot be copied in general, we have to take care of the number of copies of $\phi_{k}$ that the adversary receives. In our definition, we assume that any polynomially many copies of $\phi_{k}$ are given to the adversary. Second, we have to explicitly define what do we mean by “preimage” unlike classical OWFs, because two different quantum states can have some non-zero overlap. We introduce a QPT verification algorithm that takes $k^{\prime}$ (which is the output of the adversary) and $\phi_{k}$ as input, and outputs $\top$ or $\bot$. [^3] If $\phi_{k}$ is a pure state, we can assume without loss of generality that the verification algorithm is the following one: project |$\ket{\phi_{k}}$ onto |$\ket{\phi_{k^{\prime}}}$ and output $\top$ if the projection is successful. Otherwise, output $\bot$.
Recently, Cavalar, Goldin, Gray, Hall, Liu, and Pelecanos [${}^{+}$23], among many other results, raised the question of how short output length of OWSGs can be. They show that PRSGs with $\omega(\log\lambda)$-qubit outputs are also OWSGs, where $\lambda$ is the security parameter. [^4] A PRSG is a QPT algorithm that takes a bit string as input and outputs a quantum state whose polynomially-many copies are computationally indistinguishable from the same number of copies of Haar random states. Since it is known that OWFs imply PRSGs of arbitrary (QPT-computable) output length [BS20], their result gives OWSGs with $m(\lambda)$-qubit outputs for any QPT-computable function $m(\lambda)=\omega(\log\lambda)$ based on OWFs.555 We say that a function $m$ defined on $\mathbb{N}$ is QPT-computable if there is a QPT algorithm that computes $m(\lambda)$ on input $1^{\lambda}$ with probability $1$. Regarding a lower bound on the output length, they conjecture that there do not exist OWSGs with $O(\log\log\lambda)$-qubit outputs. However, they do not prove it and leave it open to study parameter regimes for which OWSGs do and do not exist.
1.1 Our Results
We prove that the conjecture of [${}^{+}$23] is true. Indeed, we prove a much stronger statement than their conjecture.
Theorem 1.1 ().
There do not exist OWSGs with $O(\log\lambda)$-qubit outputs.
Note that their conjecture is the impossibility of OWSGs with $O(\log\log\lambda)$-qubit outputs, but we prove the impossibility even for $O(\log\lambda)$-qubit outputs. The proof of the above theorem is quite simple. We simply consider a trivial attacker that just outputs a random input $k^{\prime}$ just ignoring the target state $\phi_{k}$. When $\phi_{k}$ is an $m$-qubit pure state, Welch bound [Wel74] (Lemma 2.4) directly implies that the above adversary has an advantage at least $2^{-m}$, which is inverse-polynomial when $m=O(\log\lambda)$. For the case of mixed state outputs, we need slightly more calculations, but we can still show that the above adversary has an inverse-polynomial advantage when $m=O(\log\lambda)$. The above theorem implies that the construction of $\omega(\log\lambda)$-output OWSGs based on OWFs from [BS20, ${}^{+}$23] is optimal in terms of the output length.
The above theorem means that when the output length is $O(\log\lambda)$, there is an adversary with advantage $1/{\mathrm{poly}}(\lambda)$. On the other hand, for OWFs with $O(\log\lambda)$-bit outputs, there is an adversary with advantage $1-1/{\mathrm{poly}}(\lambda)$ for any polynomial ${\mathrm{poly}}$ (e.g., see [Gol01, Excercise 10 of Section 2]). Thus, it is natural to ask if such an adversary with advantage $1-1/{\mathrm{poly}}(\lambda)$ exists for OWSGs with $O(\log\lambda)$-qubit outputs. First, we note that we cannot amplify the adversary’s advantage by repeating it many times in general. [^6] Note that such an amplification of the adversary’s advantage does not work even for OWFs. Indeed, if that was possible, then any weak OWF would be also a OWF. But we can easily construct a weak OWF that is not a OWF assuming the existence of OWF. Do not confuse this with amplification theorem that constructs OWFs from weak OWFs [Yao82]. The amplification theorem means that we can convert any weak OWF to a OWF, but it does not mean that any weak OWF is also a OWF. For example, suppose that an adversary given many copies of $\phi_{k}$ always outputs $k^{\prime}$ such that the verification algorithm accepts $(k^{\prime},\phi_{k})$ with probability $1/{\mathrm{poly}}(\lambda)$. Such an adversary has advantage $1/{\mathrm{poly}}(\lambda)$ against security of the OWSG, but one cannot gain anything by running it many times since it always outputs $k^{\prime}$ that has only $1/{\mathrm{poly}}(\lambda)$ acceptance probability. This leaves a possibility that we can circumvent the barrier of $O(\log\lambda)$-qubit outputs if we weaken the security requirement. To better understand the tradeoffs between the output length and security, we introduce an $\epsilon$-OWSGs, which are a parameterized variant of OWSGs where the adversary’s advantage is only required to be at most $\epsilon$, and study parameter regimes for output length for which $\epsilon$-OWSGs do and do not exist.
First, we study the regime of $\epsilon=\Theta(1)$. In this regime, we show that we can circumvent the barrier of $O(\log\lambda)$-qubit outputs.
Theorem 1.2 ().
For any constant $\epsilon>0$, there exist $\epsilon$-OWSGs with $O(\log\log\lambda)$-qubit pure state outputs assuming the existence of subexponentially secure OWFs.
We show this theorem by using quantum fingerprinting [BCWdW01]. Roughly, quantum finger printing encodes an $\ell$-bit classical string $x$ into an $O(\log(\ell)\log(\eta^{-1}))$-qubit pure state |$\ket{h_{x}}$ in such a way that $|\langle h_{x}|h_{x^{\prime}}\rangle|\leq\eta$ for any $x\neq x^{\prime}$. Our idea is to compose a OWF and quantum fingerprinting, i.e., our OWSG outputs |$\ket{h_{f(x)}}$ on input $x$ where $f$ is a OWF. When $\eta=\Theta(1)$, the length of the output is logarithmic in the output length of $f$. Thus, if we start from a OWF with ${\mathrm{polylog}}(\lambda)$-bit outputs, which exists assuming the existence of subexponentially secure OWFs, the output length of the above OWSG is $O(\log\log\lambda)$. We also show that the above construction preserves the one-wayness up to an additive constant error caused by quantum finger printing. This yields Theorem 1.2.
We show the following lower bound for the output length in this regime.
Theorem 1.3 ().
There do not exist $O(1)$-OWSGs with $(\frac{1}{2}\log\log\lambda+O(1))$-qubit outputs.
Our proof is inspired by the construction of quantum pseudorandom generators from short PRSGs [ALY23]. If their construction applies to the one-wayness, we can construct weak OWFs with $O(\log\lambda)$-bit output length, which is known to be impossible. Unfortunately, their result heavily relies on the pseudorandomness of outputs, and extending it to the OWSGs does not seem easy. Instead, we use the tomography and epsilon-nets [^7] An $\gamma$-net of $m$-qubit quantum states is a set of quantum states such that for any $m$-qubit quantum state, there exists an $\gamma$-close element in the set. We use $\gamma$ instead of $\epsilon$ to avoid confusion with the security of OWSGs. (Lemmata 2.6 and 2.7) to construct a randomized algorithm with classical outputs from the OWSGs, in which we can adapt the impossibility of weak OWFs with $O(\log\lambda)$-bit output length to break the original OWSGs. The output length is bounded by $O(\log\lambda)$ because the cardinality of $\gamma$-net in $m$-qubit space is bounded by $(C/\gamma)^{2^{2m}}$ for some constant $C$, which is polynomial for constant $\gamma$ and $m=\frac{1}{2}\log\log\lambda+O(1)$. For the OWSGs with pure state outputs, this can be improved to $\log\log\lambda+O(1).$
Next, we study the regime of $\epsilon=1-1/{\mathrm{poly}}(\lambda)$ for some polynomial ${\mathrm{poly}}$. We refer to $(1-1/{\mathrm{poly}}(\lambda))$-OWSGs as weak OWSGs following [MY22a]. We show that slightly super-constant-qubit outputs are sufficient for weak OWSGs.
Theorem 1.4 ().
For any QPT-computable function $m(\lambda)=\omega(1)$, there exist weak OWSGs with $m(\lambda)$-qubit pure state outputs assuming the existence of exponentially secure injective OWFs that map $\lambda$-bit inputs to $O(\lambda)$-bit outputs.
The main observation for proving this theorem is the following. For an angle $\theta$, let |$\ket{+_{\theta}}:=\frac{1}{\sqrt{2}}\left(\ket{0}+e^{i\theta}\ket{1}\right)$. Then we can show that for any distinct $y,y^{\prime}\in[\lambda]$, we have $|\langle+_{2\pi y/\lambda}|+_{2\pi y^{\prime}/\lambda}\rangle|\leq\sqrt{\frac{1+\cos(2\pi/\lambda)}{2}}=1-1/{\mathrm{poly}}(\lambda)$. This means that we can encode $\log\lambda$ classical bits into $1$ qubit if we allow the inner products to be as large as $1-1/{\mathrm{poly}}(\lambda)$. By repeating it, we can encode $t\log\lambda$ classical bits into $t$ qubits. By using this encoding instead of quantum fingerprinting in the construction for Theorem 1.2, we obtain weak OWSGs with $t$-qubit outputs starting from any injective OWFs with $t\log\lambda$-bit outputs. Assuming the existence of exponentially secure injective OWFs with linear expansion, there are injective OWFs with $t\log\lambda$-bit outputs for any super-constant $t$. This yields Theorem 1.4.
Finally, we show a matching lower bound for the output length of weak OWSGs.
Theorem 1.5 ().
There do not exist weak OWSGs with $O(1)$-qubit outputs.
The proof is almost the same as that of Theorem 1.3 noting that $(C/\gamma)^{2^{2m}}$ is polynomial when $\gamma$ is inverse-polynomial and $m$ is constant.
Our results are summarized in Table 1.
Table 1: Summary of the results
Security of OWSGs Upperbound of output length Lowerbound of output length
[…]
(Assuming OWFs)
constant $O(\log\log\lambda)$ [Theorem 1.2] $\frac{1}{2}\log\log\lambda+O(1)$ [Theorem 1.3]
(Assuming subexp secure OWFs)
$1-\frac{1}{{\mathrm{poly}}(\lambda)}$ $\omega(1)$ [Theorem 1.4] $O(1)$[Theorem 1.5]
(Assuming exp secure injective OWFs)
1.2 Discussion
Our results suggest that the upper and lower bounds for output length for ($\epsilon$-)OWSGs behave quite differently from those for OWFs.
  • It is easy to see that weak OWFs with $O(\log\lambda)$-bit outputs do not exist. On the other hand, assuming the existence of exponentially secure OWFs with linear expansion, we can show that there exist (strong) OWFs with $m$-bit outputs for any $m=\omega(\log\lambda)$ by a simple padding argument. This means that $O(\log\lambda)$ is the best lower bound for both OWFs and weak OWFs. On the other hand, our results suggest that the lower bounds for OWSGs and weak OWSGs are different. In particular, we show that there do not exist OWSGs with $O(\log\lambda)$-qubit outputs (Theorem 1.1), but there exist weak OWSGs with slightly super-constant-qubit outputs (Theorem 1.4), which circumvents the barrier for OWSGs.
  • For constructing OWFs with the optimal output length, i.e., those with slightly super-logarithmic-bit outputs, we need to assume exponential hardness of $\mathbf{NP}$ due to an obvious reason. On the other hand, we can construct OWSGs with slightly super-logarithmic-bit outputs only assuming polynomial hardness of OWFs [BS20, ${}^{+}$23]. Since we show that their construction is optimal in terms of output length (Theorem 1.1), this means that polynomial hardness is sufficient for achieving the optimal output length for OWSGs.
Our upper and lower bounds of the output length of OWSGs also imply those for several other primitives.
  1. 1.
    Quantum-public-key digital signatures: OWSGs are equivalent to (bounded poly-time secure) digital signatures with quantum public keys [MY22a]. Hence we have the following consequences. First, from Theorem 1.1, the length of the quantum public key of digital signatures cannot be $O(\log\lambda)$. Second, from Theorem 4.1, assuming OWFs, there exist one-time secure digital signatures with $m$-qubit quantum public keys for any $m=\omega(\log(\lambda))$. This can be regarded as a quantum advantage because it is unlikely that we can construct classical one-time secure digital signatures with similar public key length from (polynomially secure) OWFs.
  2. 2.
    Private-key quantum money: Private-key quantum money schemes with pure banknotes imply OWSGs [MY22a]. Hence, from Theorem 1.1, the length of pure quantum banknotes cannot be $O(\log\lambda)$.
Open problems.
Though we give matching upper and lower bounds for the output length of OWSGs and weak OWSGs, there is still a gap for $\Theta(1)$-OWSGs. In particular, our lower bound is $(\log\log\lambda)/2+O(1)$ (Theorem 1.3) whereas upper bound is $O(\log\log\lambda)$ (Theorem 1.2). It is an interesting open problem to study the parameter regime of constants $\epsilon$ and $c$ for which $\epsilon$-OWSGs with $c\log\log\lambda$-qubit output do and do not exist.
In this paper, we focus on $\epsilon$-OWSGs for $\epsilon\in\{{\mathsf{negl}}(\lambda),\Theta(1),1-1/{\mathrm{poly}}(\lambda)\}$, but another interesting parameter regime is $\epsilon=1/{\mathrm{poly}}(\lambda)$ for some fixed polynomial ${\mathrm{poly}}$. As for a lower bound, by using Welch bound similarly to the proof of Theorem 1.1, we would be able to show that for any polynomial ${\mathrm{poly}}$, there is a constant $c$ such that there do not exist $1/{\mathrm{poly}}(\lambda)$-OWSGs with $c\log\lambda$-qubit outputs. This still leaves a possibility that for any polynomial ${\mathrm{poly}}$, there exist $1/{\mathrm{poly}}(\lambda)$-OWSGs with $O(\log\lambda)$-qubit outputs, which circumvents the barrier for standard OWSGs. On the other hand, if we apply our construction based on quantum fingerprinting similarly to Theorem 1.2, we obtain $1/{\mathrm{poly}}(\lambda)$-OWSGs with $O((\log\lambda)(\log\log\lambda))$-qubit outputs assuming subexponentially secure OWFs. However, this is meaningless since existing works [BS20, ${}^{+}$23] already imply them based on a weaker assumption of standard OWFs. In the quantum random oracle model (QROM) [${}^{+}$11], there is quantum fingerprinting with better output length (see Remark 2.9), which leads to $1/{\mathrm{poly}}(\lambda)$-OWSGs with $O(\log\lambda)$-qubit outputs. We leave it open to construct $1/{\mathrm{poly}}(\lambda)$-OWSGs with similar parameters in the plain model (i.e., without relying on random oracles).
Another interesting question is to construct quantum OWFs using short output OWSGs. Here, a quantum OWF is a quantum (pseudo)deterministic polynomial-time algorithm whose inputs and outputs are classical bit strings and the standard one-wayness is satisfied. The construction of quantum PRGs from short output PRSGs in [ALY23] would not be applied to the case of OWSGs, because both the proofs for (pseudodeterministic) correctness and security exploit (pseudo)randomness of output states of PRSGs. Answering this question provides a better understanding of quantum cryptographic primitives with short output lengths, and simplifies the proof of Theorems 1.3 and 1.5.
1.3 Organization
We prepare definitions and technical tools in Section 2. We prove our results on lower bounds for output length of ($\epsilon$-)OWSGs (Theorems 1.1, 1.5 and 1.3) in Section 3. We prove our results on upper bounds for output length of ($\epsilon$-)OWSGs (Theorems 1.2 and 1.4) in Section 4.
2 Preliminaries
Notations.
We use standard notations of quantum computing and cryptography. We use $\lambda$ as the security parameter. $[n]$ means the set $\{1,2,...,n\}$. For a finite set $S$, $x\leftarrow S$ means that an element $x$ is sampled uniformly at random from the set $S$. ${\mathsf{negl}}$ is a negligible function, and ${\mathrm{poly}}$ is a polynomial. PPT stands for (classical) probabilistic polynomial-time and QPT stands for quantum polynomial-time. We say that a function $m$ defined on $\mathbb{N}$ is QPT-computable if there is a QPT algorithm that computes $m(\lambda)$ on input $1^{\lambda}$ with probability $1$. We stress that the running time of the algorithm can be polynomial in $\lambda$ rather than in $\log\lambda$. For an algorithm $A$, $y\leftarrow A(x)$ means that the algorithm $A$ outputs $y$ on input $x$. For quantum states $\rho,\sigma$, $\|\rho-\sigma\|_{tr}$ means their trace distance. For any pure states |$\ket{\psi},\ket{\phi}$ it is known that $\||\psi\rangle\langle\psi|-|\phi\rangle\langle\phi|\|_{tr}=\sqrt{1-|\langle\psi|\phi\rangle|^{2}}$.
Definition 2.1 (OWFs).
A function $f:\{0,1\}^{*}\to\{0,1\}^{*}$ is a (quantum-secure) OWF if it is computable in classical deterministic polynomial-time, and for any QPT adversary $\mathcal{A}$, there exists a negligible function ${\mathsf{negl}}$ such that for any $\lambda$,
$\Pr[f(x^{\prime})=f(x):x\leftarrow\{0,1\}^{\lambda},x^{\prime}\leftarrow\mathcal{A}(1^{\lambda},f(x))]\leq{\mathsf{negl}}(\lambda).$ (1)
We say that $f$ is a subexponentially secure OWF if there is a constant $10<c<$ 1 such that for any QPT adversary $\mathcal{A}$,
$\Pr[f(x^{\prime})=f(x):x\leftarrow\{0,1\}^{\lambda},x^{\prime}\leftarrow\mathcal{A}(1^{\lambda},f(x))]\leq 2^{-\lambda^{c}}$ (2)
for all sufficiently large $\lambda$.
We say that $f$ is an exponentially secure OWF if there is a constant $10<c<$ 1 such that for any QPT adversary $\mathcal{A}$,
$\Pr[f(x^{\prime})=f(x):x\leftarrow\{0,1\}^{\lambda},x^{\prime}\leftarrow\mathcal{A}(1^{\lambda},f(x))]\leq 2^{-c\lambda}$ (3)
for all sufficiently large $\lambda$.
Remark 2.2 ().
We require OWFs to be computable in classical deterministic polynomial-time following the standard definition. However, all of our positive results still follow even if we relax the requirement to only require them to be computable in QPT.
Definition 2.3 (One-way state generators (OWSGs)).
For a function $\epsilon:\mathbb{N}\rightarrow[0,1]$, an $\epsilon$-one-way state generator ($\epsilon$-OWSG) is a tuple of QPT algorithms $(\mathsf{KeyGen},\mathsf{StateGen},\mathsf{Ver})$ that work as follows:
  • $\mathsf{KeyGen}(1^{\lambda})\to k:$ It is a QPT algorithm that, on input the security parameter $\lambda$, outputs a classical key $k$.
  • $\mathsf{StateGen}(k)\to\phi_{k}:$ It is a QPT algorithm that, on input $k$, outputs a quantum state $\phi_{k}$.
  • $\mathsf{Ver}(k^{\prime},\phi_{k})\to\top/\bot:$ It is a QPT algorithm that, on input $\phi_{k}$ and a bit string $k^{\prime}$, outputs $\top$ or $\bot$.
We require the following correctness and $\epsilon$-one-wayness.
Correctness:
$\Pr[\top\leftarrow\mathsf{Ver}(k,\phi_{k}):k\leftarrow\mathsf{KeyGen}(1^{\lambda}),\phi_{k}\leftarrow\mathsf{StateGen}(k)]\geq 1-{\mathsf{negl}}(\lambda).$
$\epsilon$-one-wayness:
[…]
$\Pr[\top\leftarrow\mathsf{Ver}(k^{\prime},\phi_{k}):k\leftarrow\mathsf{KeyGen}(1^{\lambda}),\phi_{k}\leftarrow\mathsf{StateGen}(k),k^{\prime}\leftarrow\mathcal{A}(1^{\lambda},\phi_{k}^{\otimes t})]\leq\epsilon(\lambda)$
for all sufficiently large $\lambda$.
We say that it is a OWSG if it is a $\lambda^{-c}$-OWSG for all $c\in\mathbb{N}$.
We say that it is a weak OWSG if there is $c\in\mathbb{N}$ such that it is a $(1-\lambda^{-c})$-OWSG.
We say that it has $m(\lambda)$-qubit outputs if $\phi_{k}$ is an $m(\lambda)$-qubit state for all $k$ in the support of $\mathsf{KeyGen}(1^{\lambda})$.
We say that it has pure output states if $\phi_{k}$ is a pure state for all $k$ in the support of $\mathsf{KeyGen}(1^{\lambda})$.
Lemma 2.4 (Welch Bound [Wel74, BS08]).
Let $n,d,k$ be positive integers. For any family {|$\{\ket{\psi_{i}}\}_{i\in[n]}$ of (not necessarily unit) vectors in $\mathbb{C}^{d}$, it holds that
(${d+k-1\choose k}\sum_{i,j\in[n]}|\langle\psi_{i}|\psi_{j}\rangle|^{2k}\geq\left(\sum_{i\in[n]}\langle\psi_{i}|\psi_{i}\rangle^{k}\right)^{2}.$
In particular, we use the following corollary.
Corollary 2.5 ().
Let $n,m$ be positive integers. For any family {|$\{\ket{\phi_{i}}\}_{i\in[n]}$ of (normalized) $m$-qubit pure states and any distribution $\mathcal{D}$ over $[n]$, it holds that
$\mathbb{E}_{i\leftarrow\mathcal{D},j\leftarrow\mathcal{D}}|\langle\phi_{i}|\phi_{j}\rangle|^{2}\geq 2^{-m}.$
Proof.
For each $i\in[n]$, let $p_{i}:=\Pr[i\leftarrow\mathcal{D}]$ and |$\ket{\psi_{i}}:=\sqrt{p_{i}}\ket{\phi_{i}}$. Applying Lemma 2.4 for {|$\{\ket{\psi_{i}}\}_{i\in[n]}$ where $k:=1$ and $d:=2^{m}$, we obtain
$2^{m}\sum_{i,j\in[n]}p_{i}p_{j}|\langle\phi_{i}|\phi_{j}\rangle|^{2}\geq\left(\sum_{i\in[n]}p_{i}\langle\phi_{i}|\phi_{i}\rangle\right)^{2}.$
By definition, we have
$\sum_{i,j\in[n]}p_{i}p_{j}|\langle\phi_{i}|\phi_{j}\rangle|^{2}=\mathbb{E}_{i\leftarrow\mathcal{D},j\leftarrow\mathcal{D}}|\langle\phi_{i}|\phi_{j}\rangle|^{2}.$
On the other hand, since |$\ket{\phi_{i}}$ is a normalized state, we have
$\left(\sum_{i\in[n]}p_{i}\langle\phi_{i}|\phi_{i}\rangle\right)^{2}=\left(\sum_{i\in[n]}p_{i}\right)^{2}=1^{2}=1.$
Combining the above, we obtain Corollary 2.5. ∎
We use the following version of quantum state tomography given in [ALY23].
Lemma 2.6 (Quantum state tomography [ALY23, Corollary 2.16]).
Let $\delta=\delta(\lambda)\in(0,1]$ and $d=d(\lambda)\in\mathbb{N}$. There exists a tomography procedure $\mathsf{Tomography}$ that takes $t(\lambda):=144\lambda d^{4}/\delta^{2}$ copies of a $d$-dimensional mixed quantum state $\rho$, and outputs a matrix $M\in\mathbb{C}^{d\times d}$ such that
$\Pr\left[\|M-\rho\|_{tr}\leq\delta:M\leftarrow\mathsf{Tomography}(\rho^{\otimes t},\delta)\right]\geq 1-{\mathsf{negl}}(\lambda).$
Moreover, the running time of $\mathsf{Tomography}$ is polynomial in $1/\delta,d$ and $\lambda$.
Lemma 2.7 ($\gamma$-net of quantum states [AS17, Section 5]).
Let $\gamma\in(0,1]$ and $d\in\mathbb{N}$. There exist some constant $C>0$ and a set $\mathsf{Net}(d,\gamma)$ of density matrices, called an $\gamma$-net (of quantum states), of size $|\mathsf{Net}(d,\gamma)|\leq(C/\gamma)^{d^{2}}$ such that the following holds: For any $d$-dimensional mixed quantum state $\rho$, there exists $N\in\mathsf{Net}(d,\gamma)$ such that $\|N-\rho\|_{tr}\leq\gamma$. Moreover, there exists an explicit construction of $\mathsf{Net}(d,\gamma)$ in time $|\mathsf{Net}(d,\gamma)|\cdot{\mathrm{poly}}(d,1/\gamma)$. If we only consider the pure states, $|\mathsf{Net}(d,\gamma)|\leq(C/\gamma)^{d}$ holds.
Sketch of proof.
By purification, it suffices to show that there exists an explicit $\gamma$-net for pure states in dimension $d^{2}$ of size $(C/\gamma)^{d^{2}}$. This is given in [AS17, Exercise 5.22] by normalizing an (explicit) $\gamma$-net in the Boolean cube for Hamming distance in the same dimension. ∎
Lemma 2.8 (Quantum Fingerprinting [BCWdW01]).
There is a ${\mathrm{poly}}(\ell)\log\eta^{-1}$-time quantum algorithm that takes $x\in\{0,1\}^{\ell}$ as input and outputs an $O(\log(\ell)\log(\eta^{-1}))$-qubit state |$\ket{h_{x}}$ such that $|\langle h_{x}|h_{x^{\prime}}\rangle|\leq\eta$ for each $x\neq x^{\prime}$.
Proof.
Because the above statement is not explicitly written in [BCWdW01], we here provide its proof. Let $E:\{0,1\}^{\ell}\to\{0,1\}^{m}$ be an error-correcting code such that $m=c\ell$ and the Hamming distance between any two distinct code words $E(x)$ and $E(y)$ (i.e., the number of bit positions where they differ) is at least $(1-\delta)m$, where $c$ and $\delta$ are positive constants. For some specific codes, we can choose any $c>2$ and have $\delta<\frac{9}{10}+\frac{1}{15c}$ for sufficiently large $\ell$ [BCWdW01]. Let us take $c=4$. Then $\delta<\frac{9}{10}+\frac{1}{15c}=\frac{11}{12}$ for sufficiently large $\ell$. Let $r\coloneqq\lfloor\frac{\log\eta}{\log\frac{11}{12}}\rfloor+1$. For each $x\in\{0,1\}^{\ell}$, define the state
|$\ket{h_{x}}\coloneqq\left(\frac{1}{\sqrt{m}}\sum_{i=1}^{m}\ket{i}\ket{E_{i}(x)}\right)^{\otimes r},$ (4)
where $E_{i}(x)$ is the $i$th bit of $E(x)$. Then
$|\langle h_{x}|h_{x^{\prime}}\rangle|=\left(\frac{1}{m}\sum_{i=1}^{m}\langle E_{i}(x)|E_{i}(x^{\prime})\rangle\right)^{r}\leq\delta^{r}<\left(\frac{11}{12}\right)^{\frac{\log\eta}{\log\frac{11}{12}}}=\eta.$ (5)
The number $N$ of qubits of |$\ket{h_{x}}$ is
$N=r(\lceil\log(m)\rceil+1)\leq\left(\frac{\log\eta^{-1}}{\log\frac{12}{11}}+1\right)(\log(c\ell)+2)<(8\log\eta^{-1}+1)(\log\ell+4).$ (6)
Therefore we have $N=O(\log\ell\log\eta^{-1})$.
Remark 2.9 ().
The work [BCWdW01, Section 3] also shows an alternative non-constructive method for quantum fingerprinting with better output length. With some rephrasing, they show that we can take $m=\log\ell+O(\log\eta^{-1})$ in such a way that if we define |$\ket{h_{x}}:=2^{-m/2}\sum_{z\in\{0,1\}^{m}}(-1)^{H(x||z)}\ket{z}$ for a random function $H:\{0,1\}^{\ell+m}\rightarrow\{0,1\}$, then for $(1-2^{-\Omega(\ell)})$-fraction of $H$, we have $|\langle h_{x}|h_{x^{\prime}}\rangle|\leq\eta$ for all pairs of distinct $\ell$-bit strings $x\neq x^{\prime}$. Then they argue that there exists a fixed $H$ which gives a quantum fingerprinting with the desired parameter by a non-constructive argument. We observe that the above construction can be directly used as quantum fingerprinting in the quantum random oracle model (QROM) [${}^{+}$11]. This means that we have a quantum fingerprinting with output length $\log\ell+O(\log\eta^{-1})$ in the QROM, which is better than the one given in Lemma 2.8. We do not know how to achieve the same bound in the plain model (i.e., without relying on random oracles). The improvement from $O(\log(\ell)\log(\eta^{-1}))$ to $\log\ell+O(\log\eta^{-1})$ does not lead to any asymptotical improvement for output length of $\epsilon$-OWSGs in the parameter regimes which we mainly consider in this paper, but this gives an asymptotical improvement for $\epsilon=1/{\mathrm{poly}}(\lambda)$ (see the second open problem in Section 1.2).
3 Lower Bounds
In this section, we prove lower bounds for output length of ($\epsilon$-)OWSGs (Theorems 1.1, 1.5 and 1.3).
Theorem 3.1 (Restatement of Theorem 1.1).
There do not exist OWSGs with $O(\log\lambda)$-qubit outputs.
Proof.
Suppose that QPT algorithms $(\mathsf{KeyGen},\mathsf{StateGen},\mathsf{Ver})$ satisfy the syntax and correctness of OWSGs and has $m=O(\log\lambda)$-qubit outputs. We first prove the following:
$\mathbb{E}_{k,k^{\prime}}\|\phi_{k}-\phi_{k^{\prime}}\|_{tr}\leq\sqrt{1-2^{-m}}.$ (7)
We prove this below. For each $k$, we consider the diagonalization of $\phi_{k}$ to write
$\phi_{k}=\sum_{i}p_{k,i}|\phi_{k,i}\rangle\langle\phi_{k,i}|.$
Then we have
$\mathbb{E}_{k,k^{\prime}}\|\phi_{k}-\phi_{k^{\prime}}\|_{tr}$ $=\mathbb{E}_{k,k^{\prime}}\left\|\sum_{i}p_{k,i}|\phi_{k,i}\rangle\langle\phi_{k,i}|-\sum_{j}p_{k^{\prime},j}|\phi_{k^{\prime},j}\rangle\langle\phi_{k^{\prime},j}|\right\|_{tr}$ (8)
$=\mathbb{E}_{k,k^{\prime}}\left\|\sum_{i,j}p_{k,i}p_{k^{\prime},j}\left(|\phi_{k,i}\rangle\langle\phi_{k,i}|-|\phi_{k^{\prime},j}\rangle\langle\phi_{k^{\prime},j}|\right)\right\|_{tr}$ (9)
$\leq\mathbb{E}_{k,k^{\prime}}\sum_{i,j}p_{k,i}p_{k^{\prime},j}\left\||\phi_{k,i}\rangle\langle\phi_{k,i}|-|\phi_{k^{\prime},j}\rangle\langle\phi_{k^{\prime},j}|\right\|_{tr}$ (10)
$=\mathbb{E}_{k,k^{\prime}}\sum_{i,j}p_{k,i}p_{k^{\prime},j}\sqrt{1-|\langle\phi_{k,i}|\phi_{k^{\prime},j}\rangle|^{2}}$ (11)
$\leq\sqrt{1-\mathbb{E}_{k,k^{\prime}}\sum_{i,j}p_{k,i}p_{k^{\prime},j}|\langle\phi_{k,i}|\phi_{k^{\prime},j}\rangle|^{2}}$ (12)
$\leq\sqrt{1-2^{-m}},$ (13)
where Equation 10 follows from the triangle inequality for trace distance, Equation 12 follows from the concavity of the function $\sqrt{1-x}$, and Equation 13 follows from Corollary 2.5. This completes the proof of Equation 7.
Then we consider a trivial adversary $\mathcal{A}$ against its one-wayness that ignores its input (except for the security parameter) and simply outputs a fresh random key $k^{\prime}\leftarrow\mathsf{KeyGen}(1^{\lambda})$. Then we have
$\Pr[\mathcal{A}~{}\text{wins}]$ $=\mathbb{E}_{k,k^{\prime}}\Pr[\mathsf{Ver}(k^{\prime},\phi_{k})=\top]$ (14)
$\geq\mathbb{E}_{k,k^{\prime}}\left[\Pr[\mathsf{Ver}(k,\phi_{k})=\top]-\|\phi_{k}-\phi_{k^{\prime}}\|_{tr}\right]$ (15)
$\geq 1-{\mathsf{negl}}(\lambda)-\sqrt{1-2^{-m}},$ (16)
where Equation 16 follows from the correctness of $(\mathsf{KeyGen},\mathsf{StateGen},\mathsf{Ver})$ and Equation 7. Note that we have $\sqrt{1-x}<1-x/4$ for all $0\leq x\leq\frac{1}{2}$ and $\sqrt{1-x}\leq\sqrt{\frac{1}{2}}$ for all $\frac{1}{2}\leq x\leq 1$. Therefore we have
$\Pr[\mathcal{A}~{}\text{wins}]\geq\min\left\{2^{-m-2},1-\sqrt{\frac{1}{2}}\right\}-{\mathsf{negl}}(\lambda).$ (17)
Since $m=O(\log\lambda)$, the RHS is non-negligible and thus $\mathcal{A}$ succeeds in breaking one-wayness. ∎
Remark 3.2 ().
In the above proof, $\mathcal{A}$ does not use the target state at all. This means that even zero-copy OWSGs (where the adversary is given no copy of the target state) with $O(\log\lambda)$-qubit outputs do not exist.
For proving Theorems 1.5 and 1.3, we prove the following theorem.
Theorem 3.3 ().
Let $\Delta=\Delta(\lambda)>0$ and $m=m(\lambda)\in\mathbb{N}$ and $C$ be a constant taken from Lemma 2.7. If $(6C/\Delta)^{2^{2m}}={\mathrm{poly}}(\lambda)$, there do not exist $(1-\Delta)$-OWSGs with $m$-qubit outputs.
Proof.
Suppose that QPT algorithms $(\mathsf{KeyGen},\mathsf{StateGen},\mathsf{Ver})$ satisfy the syntax and correctness of $(1-\Delta)$-OWSGs with $m$-qubit outputs. Let $\gamma=\Delta/6$, $d=2^{m}$, $t=144\lambda d^{4}/\gamma^{2}$, and $T=\frac{(C/\gamma)^{d^{2}}\log\lambda}{\gamma}$. We consider the following adversary $\mathcal{A}(1^{\lambda},\phi_{k}^{\otimes t})$:
  1. 1.
    Construct an $\gamma$-net $\mathsf{Net}(d,\gamma)$.
  2. 2.
    Given $\phi_{k}^{\otimes t}$, run $M\leftarrow\mathsf{Tomography}(\phi_{k}^{\otimes t},\gamma)$.
  3. 3.
    Construct a set $\mathcal{N}=\{N\in\mathsf{Net}(d,\gamma):\|N-M\|_{tr}\leq\gamma\}.$
  4. 4.
    For each $i\in[T]$, do:
    1. (a)
      Sample $k_{i}\leftarrow\mathsf{KeyGen}(1^{\lambda})$ and run $M_{i}\leftarrow\mathsf{Tomography}(\phi_{k_{i}}^{\otimes t},\gamma).$
    2. (b)
      If there exists $N\in\mathcal{N}$ such that $\|M_{i}-N\|_{tr}\leq\gamma$, output $k^{\prime}:=k_{i}$ and terminate.
  5. 5.
    Output $\bot$.
Because of Lemmata 2.6 and 2.7, the time complexity of this algorithm is bounded by a polynomial in $1/\gamma,d,\lambda,T$ and $|\mathsf{Net}(d,\gamma)$. By the condition, $|\mathsf{Net}(d,\gamma)|\leq(C/\gamma)^{d^{2}}=(6C/\Delta)^{2^{2m}}={\mathrm{poly}}(\lambda)$ and this implies that $1/\Delta={\mathrm{poly}}(\lambda)$ and $d=2^{m}={\mathrm{poly}}(\lambda)$. This proves that $\mathcal{A}$ is a QPT algorithm.
To prove the correctness, we first argue that $\mathcal{A}$ outputs $\bot$ with a small probability. Let $\epsilon_{bad}:=\frac{\gamma}{|\mathsf{Net}(d,\gamma)|}$. Let $\mathcal{B}$ be the set of elements $N\in\mathsf{Net}(d,\gamma)$ such that the following inequality holds:
$\Pr\left[\|N-M^{\prime}\|_{tr}\leq\gamma:k^{\prime}\leftarrow\mathsf{KeyGen}(1^{\lambda}),M^{\prime}\leftarrow\mathsf{Tomography}(\phi_{k^{\prime}}^{\otimes t},\gamma)\right]\leq\epsilon_{bad}.$ (18)
Let $\sf{bad}$ be the event that $\mathcal{N}\cap\mathcal{B}\neq\emptyset$. It holds that
$\Pr[\sf{bad}]=\Pr[\mathcal{N}\cap\mathcal{B}\neq\emptyset]\leq\sum_{N\in\mathcal{B}}\Pr[N\in\mathcal{N}]\leq|\mathcal{B}|\cdot\epsilon_{bad}\leq\gamma,$ (19)
where we have used $|\mathcal{B}|\leq|\mathsf{Net}(d,\gamma)|$ at the last inequality.
Suppose that $\sf{bad}$ does not occur. Then, since $\mathcal{N}$ is not empty, there exists an element $N\in\mathcal{N}$ such that
$\Pr\left[\|N-M^{\prime}\|_{tr}\leq\gamma:k^{\prime}\leftarrow\mathsf{KeyGen}(1^{\lambda}),M^{\prime}\leftarrow\mathsf{Tomography}(\phi_{k^{\prime}}^{\otimes t},\gamma)\right]>\epsilon_{bad},$ (20)
which means that each iteration of the fourth step terminates with probability at least $\frac{\gamma}{|\mathsf{Net}(d,\gamma)|}$. Then we have
$\Pr\left[k^{\prime}\neq\bot:k\leftarrow\mathsf{KeyGen}(1^{\lambda}),k^{\prime}\leftarrow\mathcal{A}(1^{\lambda},\phi_{k}^{\otimes t})\right]$ (21)
$\geq\Pr\left[k^{\prime}\neq\bot:k\leftarrow\mathsf{KeyGen}(1^{\lambda}),k^{\prime}\leftarrow\mathcal{A}(1^{\lambda},\phi_{k}^{\otimes t})|\lnot{\sf{bad}}\right]\cdot\Pr[\lnot{\sf{bad}}]$ (22)
$\geq\left(1-\left(1-\epsilon_{bad}\right)^{T}\right)\cdot(1-\gamma)\geq 1-\gamma-{\mathsf{negl}}(\lambda),$ (23)
where the first inequality in Equation 23 holds because the iterations are independent, and the last inequality holds because of the choice of $T\geq{\log\lambda}/{\epsilon_{bad}}$.
Finally, we prove that the output of $\mathcal{A}$ is accepted by $\mathsf{Ver}$ with high probability whenever it is not $\bot$. If $\mathcal{A}$ outputs $k_{i}$, $\|\phi_{k}-\phi_{k_{i}}\|_{tr}\leq 4\gamma$ with probability $1-{\mathsf{negl}}(\lambda)$. By the triangular inequality, we have
$\|\phi_{k}-\phi_{k_{i}}\|_{tr}\leq\|\phi_{k}-M\|_{tr}+\|M-N\|_{tr}+\|N-M_{i}\|_{tr}+\|M_{i}-\phi_{k_{i}}\|_{tr},$ (24)
all of which are bounded by $\gamma$ with probability $1-{\mathsf{negl}}(\lambda)$ because of Lemma 2.6 and the definition of $\mathsf{Net}(d,\gamma)$. This implies that
$\Pr[\top\leftarrow\mathsf{Ver}(k_{i},\phi_{k})]\geq\Pr[\top\leftarrow\mathsf{Ver}(k,\phi_{k})]-\|\phi_{k}-\phi_{k_{i}}\|_{tr}\geq 1-4\gamma-{\mathsf{negl}}(\lambda)$ (25)
holds with probability $1-{\mathsf{negl}}(\lambda)$, where the last inequality follows from the correctness of $(\mathsf{KeyGen},\mathsf{StateGen},\mathsf{Ver})$. Using $\gamma=\Delta/6$, the advantage of $\mathcal{A}$ is bounded as follows:
[…]
$\geq(1-4\gamma-{\mathsf{negl}}(\lambda))\cdot(1-\gamma-{\mathsf{negl}}(\lambda))\cdot(1-{\mathsf{negl}}(\lambda)$ (27)
$\geq 1-5\gamma-{\mathsf{negl}}(\lambda)\geq 1-\Delta$ (28)
where the first term of Equation 27 is from Equation 25, the second term is from Equation 23, and the last term is the probability that Equation 25 holds. Therefore, $\mathcal{A}$ is a QPT algorithm breaking $(1-\Delta)$-OWSG of $(\mathsf{KeyGen},\mathsf{StateGen},\mathsf{Ver})$. ∎
Theorems 1.5 and 1.3 follow from Theorem 3.3 as immediate corollaries.
Corollary 3.4 (Restatement of Theorem 1.3).
There do not exist $\epsilon$-OWSGs with $m$-qubit outputs for any constant $\epsilon>0$ and $m=\frac{1}{2}\log\log\lambda+O(1)$.
Proof.
If the output length $m$ is $m=\frac{1}{2}\log\log\lambda+O(1)$ so that $2^{2m}=O(\log\lambda),$ then for any constant $\epsilon>0$, $(6C/(1-\epsilon))^{2^{2m}}=(6C/(1-\epsilon))^{O(\log\lambda)}={\mathrm{poly}}(\lambda)$ and we can apply Theorem 3.3. ∎
Remark 3.5 ().
For the pure case, we can prove that there do not exist $\epsilon$-OWSGs with $(\log\log\lambda+O(1))$-qubit outputs for any constant $\epsilon>0$ using the pure case of Lemma 2.7.
Corollary 3.6 (Restatement of Theorem 1.5).
There do not exist weak OWSGs with any constant output length.
Proof.
Consider the security of weak OWSGs with the constant output length $m$, which is $(1-\epsilon)$-secure for some inverse polynomial $\epsilon$. Since $m$ is constant, $(6C/\epsilon)^{2^{2m}}={\mathrm{poly}}(\lambda)$, thus Theorem 3.3 applies. ∎
4 Upper Bounds
In this section, we prove upper bounds for output length of ($\epsilon$-)OWSGs (Theorems 1.2 and 1.4).
We first observe that the combination of previous results already gives OWSGs with slightly super-logarithmic-qubit outputs.
Theorem 4.1 ([BS20, ${}^{+}$23]).
If there exist OWFs, then there exist OWSGs with $m(\lambda)$-qubit pure state outputs for any QPT-computable function $m(\lambda)=\omega(\log\lambda)$.
Proof.
According to [BS20], if there exist OWFs, then there exist PRSGs with input length $\lambda$ and any output length $m$. From [${}^{+}$23, Theorem 3.4], these PRSGs with $m=\omega(\log\lambda)$ are also OWSGs. ∎
For proving Theorem 1.2, we prove the following theorem.
Theorem 4.2 ().
If there exist OWFs that map $\lambda$-bit inputs to $\ell(\lambda)$-bit outputs, then there exist $(\eta(\lambda)^{2}+{\mathsf{negl}}(\lambda))$-OWSGs with $O(\log(\ell(\lambda))\log(\eta(\lambda)^{-1}))$-qubit pure state outputs for any QPT-computable function $\eta(\lambda)$.
Proof.
For notational simplicity, we simply write $\ell$ and $\eta$ to mean $\ell(\lambda)$ and $\eta(\lambda)$, respecitively. Let $f$ be a OWF that maps $\lambda$-bit inputs to $\ell$-bit outputs. We construct an $(\eta^{2}+{\mathsf{negl}}(\lambda))$-OWSG as follows:
  • $\mathsf{KeyGen}(1^{\lambda})\to k:$ Sample $k\leftarrow\{0,1\}^{\lambda}$.
  • $\mathsf{StateGen}(k)\to\ket{\phi_{k}}:$ Output |$\ket{\phi_{k}}\coloneqq\ket{h_{f(k)}}$, where |$\ket{h_{f(k)}}$ is the quantum fingerprinting state of $f(k)$ given in Lemma 2.8.
  • $\mathsf{Ver}(k^{\prime},\ket{\phi_{k}})\to\top/\bot:$ Project |$\ket{\phi_{k}}$ onto |$\ket{\phi_{k^{\prime}}}$. If the projection is successful, output $\top$. Otherwise, output $\bot$.
Let $\mathcal{A}$ be a QPT adversary of this OWSG that receives $t$ copies of |$\ket{\phi_{k}}$. Then, as we will show below, there exists a QPT adversary $\mathcal{B}$ of the OWF $f$ that satisfies [^9] We say that $\mathcal{A}$ wins if $\mathsf{Ver}(k^{\prime},\ket{\phi_{k}})\rightarrow\top$ where $k\leftarrow\mathsf{KeyGen}(1^{\lambda})$ and $k^{\prime}\leftarrow\mathcal{A}(1^{\lambda},\ket{\phi_{k}}^{\otimes t})$. Similarly, we say that $\mathcal{B}$ wins if $f(x^{\prime})=f(x)$ where $x\leftarrow\{0,1\}^{\lambda}$ and $x^{\prime}\leftarrow\mathcal{B}(1^{\lambda},f(x))$.
$\Pr[\mathcal{A}~{}\mbox{wins}]\leq\Pr[\mathcal{B}~{}\mbox{wins}]+\eta^{2}.$ (29)
Because $f$ is a OWF, we have
$\Pr[\mathcal{A}~{}\mbox{wins}]\leq{\mathsf{negl}}(\lambda)+\eta^{2},$ (30)
which shows that the OWSG is $(\eta^{2}+{\mathsf{negl}}(\lambda))$-secure. From Lemma 2.8, the output length of the OWSG is at most $O(\log(\ell)\log(\eta^{-1}))$.
Now we show Equation 29. Let $\mathcal{A}$ be a QPT adversary of the OWSG. From this $\mathcal{A}$, we construct an adversary $\mathcal{B}$ of the OWF $f$ as follows:
  1. 1.
    Given $(1^{\lambda},f(k))$, run $k^{\prime}\leftarrow\mathcal{A}(1^{\lambda},\ket{h_{f(k)}}^{\otimes t})$.
  2. 2.
    Output $k^{\prime}$.
From Lemma 2.8, $|\langle h_{f(k)}|h_{f(k^{\prime})}\rangle|\leq\eta$ if $f(k)\neq f(k^{\prime})$. Therefore we have
$\Pr[\mathcal{A}~{}\mbox{wins}]$ $=\frac{1}{2^{\lambda}}\sum_{k}\sum_{k^{\prime}}\Pr[k^{\prime}\leftarrow\mathcal{A}(1^{\lambda},\ket{h_{f(k)}}^{\otimes t})]|\langle h_{f(k)}|h_{f(k^{\prime})}\rangle|^{2}$ (31)
$=\frac{1}{2^{\lambda}}\sum_{k}\sum_{k^{\prime}:f(k^{\prime})=f(k)}\Pr[k^{\prime}\leftarrow\mathcal{A}(1^{\lambda},\ket{h_{f(k)}}^{\otimes t})]|\langle h_{f(k)}|h_{f(k^{\prime})}\rangle|^{2}$ (32)
$+\frac{1}{2^{\lambda}}\sum_{k}\sum_{k^{\prime}:f(k^{\prime})\neq f(k)}\Pr[k^{\prime}\leftarrow\mathcal{A}(1^{\lambda},\ket{h_{f(k)}}^{\otimes t})]|\langle h_{f(k)}|h_{f(k^{\prime})}\rangle|^{2}$ (33)
$\leq\frac{1}{2^{\lambda}}\sum_{k}\sum_{k^{\prime}:f(k^{\prime})=f(k)}\Pr[k^{\prime}\leftarrow\mathcal{A}(1^{\lambda},\ket{h_{f(k)}}^{\otimes t})]$ (34)
$+\frac{1}{2^{\lambda}}\sum_{k}\sum_{k^{\prime}:f(k^{\prime})\neq f(k)}\Pr[k^{\prime}\leftarrow\mathcal{A}(1^{\lambda},\ket{h_{f(k)}}^{\otimes t})]\eta^{2}$ (35)
$\leq\frac{1}{2^{\lambda}}\sum_{k}\sum_{k^{\prime}:f(k^{\prime})=f(k)}\Pr[k^{\prime}\leftarrow\mathcal{A}(1^{\lambda},\ket{h_{f(k)}}^{\otimes t})]+\eta^{2}$ (36)
$\leq\Pr[\mathcal{B}~{}\mbox{wins}]+\eta^{2}.$ (37)
Theorem 1.2 follows from Theorem 4.2 as an immediate corollary.
Corollary 4.3 (Restatement of Theorem 1.2).
For any constant $\epsilon>0$, there exist $\epsilon$-OWSGs with $O(\log\log\lambda)$-qubit pure state outputs assuming the existence of subexponentially secure OWFs.
Proof.
Without loss of generality, we can assume that a subexponentially secure OWF is length-preserving. i.e., it maps $\lambda$-bit inputs to $\lambda$-bit outputs. Let $f$ be a length-preserving subexponentially secure OWFs. Since $f$ is subexponentially secure there is a constant $10<c<$ 1 such that for any QPT adversary $\mathcal{A}$,
$\Pr[f(x^{\prime})=f(x):x\leftarrow\{0,1\}^{\lambda},x^{\prime}\leftarrow\mathcal{A}(1^{\lambda},f(x))]\leq 2^{-\lambda^{c}}$ (38)
for all sufficiently large $\lambda$. Then we construct a OWF $g$ that maps $\lambda$-bit inputs to $\ell(\lambda):=\lfloor(\log\lambda)^{2c^{-1}}\rfloor$-bit outputs as follows: for any $x\in\{0,1\}^{\lambda}$, $g(x):=f(x_{[\ell]})$ where $x_{[\ell]}$ is the first $\ell$ bits of $x$. By a straightforward reduction, for any QPT adversary $\mathcal{A}$, we have
$\Pr[g(x^{\prime})=g(x):x\leftarrow\{0,1\}^{\lambda},x^{\prime}\leftarrow\mathcal{A}(1^{\lambda},g(x))]\leq 2^{-\ell^{c}}\leq 2^{-(\log\lambda)^{2}}={\mathsf{negl}}(\lambda).$ (39)
Thus, $g$ is a OWF that maps $\lambda$-bit inputs to $\ell(\lambda)=\lfloor(\log\lambda)^{2c^{-1}}\rfloor$-bit outputs. By applying Theorem 4.2 for $\eta:=(\epsilon/2)^{1/2}$, there is a $(\epsilon/2+{\mathsf{negl}}(\lambda))$-OWSG with $O(\log\ell)=O(\log\log\lambda)$-qubit pure state outputs. Since $\epsilon/2+{\mathsf{negl}}(\lambda)<\epsilon$ for sufficiently large $\lambda$, this implies Corollary 4.3. ∎
For proving Theorem 1.4, we prove the following theorem.
Theorem 4.4 ().
If there exist injective OWFs that map $\lambda$-bit inputs to $\ell(\lambda)$-bit outputs, then there exist weak OWSGs with $\lceil\frac{\ell(\lambda)}{\log\lambda}\rceil$-qubit outputs.
Proof.
We simply write $\ell$ to mean $\ell(\lambda)$ for notational simplicity. Let $f$ be an injective OWF that maps $\lambda$-bit inputs to $\ell$-bit outputs. Define $t:=\lceil\frac{\ell}{\log\lambda}\rceil$, and we embed $\{0,1\}^{\ell}$ into $[\lambda]^{t}$. Then we can regard $f$ as a function that maps $\lambda$-bit strings to an element of $[\lambda]^{t}$. For $\theta$, let
|$\ket{+_{\theta}}:=\frac{1}{\sqrt{2}}(\ket{0}+e^{i\theta}\ket{1}).$
We construct a OWSG as follows:
  • $\mathsf{KeyGen}(1^{\lambda})\to k:$ Sample $k\leftarrow\{0,1\}^{\lambda}$.
  • $\mathsf{StateGen}(k)\to\ket{\phi_{k}}:$ Compute $f(k)=(y_{1},y_{2},...,y_{t})\in[\lambda]^{t}$ and output |$\ket{\phi_{k}}\coloneqq\bigotimes_{j\in[t]}\ket{+_{2\pi y_{j}/\lambda}}$.
  • $\mathsf{Ver}(k^{\prime},\ket{\phi_{k}})\to\top/\bot:$ Project |$\ket{\phi_{k}}$ onto |$\ket{\phi_{k^{\prime}}}$. If the projection is successful, output $\top$. Otherwise, output $\bot$.
For any $y\neq y^{\prime}$, we have $|\langle+_{2\pi y/\lambda}|+_{2\pi y^{\prime}/\lambda}\rangle|\leq\sqrt{\frac{1+\cos(2\pi/\lambda)}{2}}$. This is because
$|\langle+_{2\pi y/\lambda}|+_{2\pi y^{\prime}/\lambda}\rangle|$ $=\left|\frac{1+e^{i2\pi(y-y^{\prime})/\lambda}}{2}\right|$ (40)
$=\left|\frac{1+\cos(2\pi(y-y^{\prime})/\lambda)+i\sin(2\pi(y-y^{\prime}))}{2}\right|$ (41)
$=\sqrt{\left(\frac{1+\cos(2\pi(y-y^{\prime})/\lambda)}{2}\right)^{2}+\frac{\sin^{2}(2\pi(y-y^{\prime}))}{4}}$ (42)
$\leq\sqrt{\frac{1+\cos(2\pi(y-y^{\prime})/\lambda)}{2}}$ (43)
$\leq\sqrt{\frac{1+\cos(2\pi/\lambda)}{2}}.$ (44)
Moreover, since $f$ is injective, for any $k\neq k^{\prime}$, $f(k)$ and $f(k^{\prime})$ differ on at least one coordinate. This implies that for any $k\neq k^{\prime}$, we have $|\langle\phi_{k}|\phi_{k^{\prime}}\rangle|\leq\sqrt{\frac{1+\cos(2\pi/\lambda)}{2}}$. By a similar reduction to that in the proof of Theorem 4.2, we can see that the above OWSG satisfies $\left(\frac{1+\cos(2\pi/\lambda)}{2}+{\mathsf{negl}}(\lambda)\right)$-one-wayness. Since $\cos\theta=1-\theta^{2}/2+O(\theta^{4})$, we have $\cos(2\pi/\lambda)\leq 1-(\pi/\lambda)^{2}$ for sufficiently large $\lambda$. Thus, $\frac{1+\cos(2\pi/\lambda)}{2}=1-\frac{1}{{\mathrm{poly}}(\lambda)}$ and thus the above is a weak OWSG. ∎
Theorem 1.4 follows from Theorem 4.4 as a immediate corollary.
Corollary 4.5 (Restatement of Theorem 1.4).
For any QPT-computable function $m(\lambda)=\omega(1)$, there exist weak OWSGs with $m(\lambda)$-qubit pure state outputs assuming the existence of exponentially secure injective OWFs that map $\lambda$-bit inputs to $O(\lambda)$-bit outputs.
Proof.
Suppose that $f$ is an exponentially secure injective OWF that maps $\lambda$-bit inputs to $k\lambda$-bit outputs for some constant $k\geq 1$. Since it is exponentially secure, there is a constant $10<c<$ 1 such that for any QPT adversary $\mathcal{A}$,
$\Pr[f(x^{\prime})=f(x):x\leftarrow\{0,1\}^{\lambda},x^{\prime}\leftarrow\mathcal{A}(1^{\lambda},f(x))]\leq 2^{-c\lambda}$ (45)
for all sufficiently large $\lambda$. Then we construct a OWF $g$ that maps $\lambda$-bit inputs to $\ell(\lambda):=\lfloor m(\lambda)\log\lambda\rfloor$-bit outputs as follows: for any $x\in\{0,1\}^{\lambda}$, $g(x):=f(x_{[\ell]})$ where $x_{[\ell]}$ is the first $\ell$ bits of $x$. By a straightforward reduction, for any QPT adversary $\mathcal{A}$, we have
$\Pr[g(x^{\prime})=g(x):x\leftarrow\{0,1\}^{\lambda},x^{\prime}\leftarrow\mathcal{A}(1^{\lambda},g(x))]\leq 2^{-c\ell}\leq 2^{c}\cdot\lambda^{-cm(\lambda)}={\mathsf{negl}}(\lambda),$ (46)
where the final equality follows from $m(\lambda)=\omega(1)$. Thus, $g$ is a OWF that maps $\lambda$-bit inputs to $\ell(\lambda)=\lfloor m(\lambda)\log\lambda\rfloor$-bit outputs. By applying Theorem 4.4, there is a weak OWSG with $\lceil\frac{\ell(\lambda)}{\log\lambda}\rceil$-qubit outputs. By the definition of $\ell(\lambda)$, we have $\lceil\frac{\ell(\lambda)}{\log\lambda}\rceil\leq m(\lambda)$. Since we can arbitrarily increase the output length of weak OWSGs by a trivial padding, this implies Corollary 4.5. ∎
Acknowledgements. This work was initiated while the authors were joining NII Shonan Meeting No.198 “New Directions in Provable Quantum Advantages”. MH is supported by a KIAS Individual Grant QP089801. TM is supported by JST CREST JPMJCR23I3, JST Moonshot R&D JPMJMS2061-5-1-1, JST FOREST, MEXT QLEAP, the Grant-in Aid for Transformative Research Areas (A) 21H05183, and the Grant-in-Aid for Scientific Research (A) No.22H00522.