### v3

# A Note on Output Length of One-Way State Generators and EFIs

Abstract
We study the output length of one-way state generators (OWSGs), their weaker variants, and EFIs.
  • Standard OWSGs.
    Recently, Cavalar et al. (arXiv:2312.08363) give OWSGs with $m$-qubit outputs for any $m=\omega(\log\lambda)$, where $\lambda$ is the security parameter, and conjecture that there do not exist OWSGs with $O(\log\log\lambda)$-qubit outputs. We prove their conjecture in a stronger manner by showing that there do not exist OWSGs with $O(\log\lambda)$-qubit outputs. This means that their construction is optimal in terms of output length.
  • Inverse-polynomial-advantage OWSGs.
    Let $\epsilon$-OWSGs be a parameterized variant of OWSGs where a quantum polynomial-time adversary’s advantage is at most $\epsilon$. For any constant $c\in\mathbb{N}$, we construct $\lambda^{-c}$-OWSGs with $((c+1)\log\lambda+O(1))$-qubit outputs assuming the existence of OWFs. We show that this is almost tight by proving that there do not exist $\lambda^{-c}$-OWSGs with at most $(c\log\lambda-2)$-qubit outputs.
  • Constant-advantage OWSGs.
    For any constant $\epsilon>0$, we construct $\epsilon$-OWSGs with $O(\log\log\lambda)$-qubit outputs assuming the existence of subexponentially secure OWFs. We show that this is almost tight by proving that there do not exist $O(1)$-OWSGs with $((\log\log\lambda)/2+O(1))$-qubit outputs.
  • Weak OWSGs.
    We refer to $(1-1/{\mathrm{poly}}(\lambda))$-OWSGs as weak OWSGs. We construct weak OWSGs with $m$-qubit outputs for any $m=\omega(1)$ assuming the existence of exponentially secure OWFs with linear expansion. We show that this is tight by proving that there do not exist weak OWSGs with $O(1)$-qubit outputs.
  • EFIs.
    We show that there do not exist $O(\log\lambda)$-qubit EFIs. We show that this is tight by proving that there exist $m(\lambda)$-qubit EFIs for any $m(\lambda)=\omega(\log\lambda)$ assuming the existence of exponentially secure pseudorandom generators.
1 Introduction
In classical cryptography, the existence of one-way functions (OWFs) is the minimum assumption, because many primitives (such as secret-key encryption (SKE), commitments, zero-knowledge, pseudorandom generators (PRGs), pseudorandom functions (PRFs), and digital signatures) are equivalent to OWFs, and almost all cryptographic primitives imply OWFs [LR86, IL89, ILL89]. On the other hand, in quantum cryptography where quantum computing and quantum communications are available, recent studies suggest that OWFs would not be the minimum assumption, and many new replacements of OWFs have been introduced. For example, pseudorandom state generators (PRSGs) [JLS18], one-way state generators (OWSGs) [MY22b, MY22a], and EFIs [BCQ23] are quantum analogue of PRGs, OWFs, and EFIDs [Gol90], respectively. They seem to be weaker than OWFs [Kre21, KQST23, LMW23], yet can be used to construct several useful primitives such as private-key quantum money, SKE, commitments, multiparty computations, and digital signatures [MY22b, AQY22, BCQ23, Yan22].
OWSGs are a quantum analog of OWFs.111 There are a pure state output version [MY22b] and mixed state output version [MY22a] of OWSGs. We consider OWSGs with mixed state outputs as a default definition, but all of our positive results actually have pure state outputs. We adopt the mixed state version to make our negative results stronger. A OWSG is a quantum polynomial-time (QPT) algorithm that takes a bit string $k$ as input and outputs a quantum state $\phi_{k}$. The security is, roughly speaking, that no QPT adversary can find “preimage” of $\phi_{k}$. There are two differences from classical OWFs. First, because quantum states cannot be copied in general, we have to take care of the number of copies of $\phi_{k}$ that the adversary receives. In our definition, we assume that any polynomially many copies of $\phi_{k}$ are given to the adversary. Second, we have to explicitly define what do we mean by “preimage” unlike classical OWFs, because two different quantum states can have some non-zero overlap. We introduce a QPT verification algorithm that takes $k^{\prime}$ (which is the output of the adversary) and $\phi_{k}$ as input, and outputs $\top$ or $\bot$. [^2] If $\phi_{k}$ is a pure state, we can assume without loss of generality that the verification algorithm is the following one: project |$\ket{\phi_{k}}$ onto |$\ket{\phi_{k^{\prime}}}$ and output $\top$ if the projection is successful. Otherwise, output $\bot$.
Recently, Cavalar, Goldin, Gray, Hall, Liu, and Pelecanos [CGG+23], among many other results, raised the question of how short output length of OWSGs can be. They show that PRSGs with $\omega(\log\lambda)$-qubit outputs are also OWSGs, where $\lambda$ is the security parameter. [^3] A PRSG is a QPT algorithm that takes a bit string as input and outputs a quantum state whose polynomially-many copies are computationally indistinguishable from the same number of copies of Haar random states. Since it is known that OWFs imply PRSGs of arbitrary (QPT-computable) output length [BS20], their result gives OWSGs with $m(\lambda)$-qubit outputs for any QPT-computable function $m(\lambda)=\omega(\log\lambda)$ based on OWFs.444 We say that a function $m$ defined on $\mathbb{N}$ is QPT-computable if there is a QPT algorithm that computes $m(\lambda)$ on input $1^{\lambda}$ with probability $1$. Regarding a lower bound on the output length, they conjecture that there do not exist OWSGs with $O(\log\log\lambda)$-qubit outputs. However, they do not prove it and leave it open to study parameter regimes for which OWSGs do and do not exist.
An EFI is a pair of efficiently generatable states that are statistically far but computationally indistinguishable. While relationships between EFIs and other primitives have been actively studied [MY22a, BCQ23], to our knowledge, there is no existing work that studies how many qubits are needed for EFIs.
1.1 Our Results
Lower Bounds for ($\epsilon$-)OWSGs.
We prove that the conjecture of [CGG+23] about the lower bound for the output length of OWSGs is true. Indeed, we prove a much stronger statement than their conjecture.
Theorem 1.1.
There do not exist OWSGs with $O(\log\lambda)$-qubit outputs.
Note that their conjecture is the impossibility of OWSGs with $O(\log\log\lambda)$-qubit outputs, but we prove the impossibility even for $O(\log\lambda)$-qubit outputs. The proof of the above theorem is quite simple. We simply consider a trivial attacker that just outputs a random input $k^{\prime}$ just ignoring the target state $\phi_{k}$. When $\phi_{k}$ is an $m$-qubit pure state, Welch bound [Wel74] (Lemma 2.8) directly implies that the above adversary has an advantage at least $2^{-m}$, which is inverse-polynomial when $m=O(\log\lambda)$. For the case of mixed state outputs, we need slightly more calculations, but we can still show that the above adversary has an inverse-polynomial advantage when $m=O(\log\lambda)$.
The above theorem means that when the output length is $O(\log\lambda)$, there is an adversary with advantage $1/{\mathrm{poly}}(\lambda)$. On the other hand, for OWFs with $O(\log\lambda)$-bit outputs, there is an adversary with advantage $1-1/{\mathrm{poly}}(\lambda)$ for any polynomial ${\mathrm{poly}}$ (e.g., see [Gol01, Excercise 10 of Section 2]). Thus, it is natural to ask if such an adversary with advantage $1-1/{\mathrm{poly}}(\lambda)$ exists for OWSGs with $O(\log\lambda)$-qubit outputs. First, we note that we cannot amplify the adversary’s advantage by repeating it many times in general. [^5] Note that such an amplification of the adversary’s advantage does not work even for OWFs. Indeed, if that was possible, then any weak OWF would be also a OWF. But we can easily construct a weak OWF that is not a OWF assuming the existence of OWF. Do not confuse this with the amplification theorem that constructs OWFs from weak OWFs [Yao82]. The amplification theorem means that we can convert any weak OWF to a OWF, but it does not mean that any weak OWF is also a OWF. For example, suppose that an adversary is given many copies of $\phi_{k}$ always outputs $k^{\prime}$ such that the verification algorithm accepts $(k^{\prime},\phi_{k})$ with probability $1/{\mathrm{poly}}(\lambda)$. Such an adversary has an advantage $1/{\mathrm{poly}}(\lambda)$ against the security of the OWSG, but one cannot gain anything by running it many times since it always outputs $k^{\prime}$ that has only $1/{\mathrm{poly}}(\lambda)$ acceptance probability. This leaves a possibility that we can circumvent the barrier of $O(\log\lambda)$-qubit outputs if we weaken the security requirement. To better understand the tradeoffs between the output length and security, we introduce an $\epsilon$-OWSGs, which are a parameterized variant of OWSGs where the adversary’s advantage is only required to be at most $\epsilon$, and study parameter regimes for output length for which $\epsilon$-OWSGs do and do not exist.
First, we study the regime of $\epsilon=\lambda^{-\Theta(1)}$. We show that the proof of Theorem 1.1 naturally extends to this setting to obtain the following theorem.
Theorem 1.2.
For any constant $c\in\mathbb{N}$, there do not exist $\lambda^{-c}$-OWSGs of which output length is at most $c\log\lambda-2$.
Next, we study the regime of $\epsilon=O(1)$. We show the following lower bound.
Theorem 1.3.
There do not exist $O(1)$-OWSGs with $(\frac{1}{2}\log\log\lambda+O(1))$-qubit outputs.
In this regime, the idea for the proofs of Theorems 1.1 and 1.2 does not work since Welch bound only gives an adversary with advantage $1/{\mathrm{poly}}(\lambda)$. Thus, we rely on a different idea. Our proof is inspired by the construction of quantum pseudorandom generators from short PRSGs [ALY23]. If their construction applies to the one-wayness, we can construct weak OWFs with $O(\log\lambda)$-bit output length, which is known to be impossible. Unfortunately, their result heavily relies on the pseudorandomness of outputs, and extending it to the OWSGs does not seem easy. Instead, we use the tomography and epsilon-nets [^6] An $\gamma$-net of $m$-qubit quantum states is a set of quantum states such that for any $m$-qubit quantum state, there exists an $\gamma$-close element in the set. We use $\gamma$ instead of $\epsilon$ to avoid confusion with the security of OWSGs. (Lemmata 2.10 and 2.12) to construct a randomized algorithm with classical outputs from the OWSGs, in which we can adapt the impossibility of weak OWFs with $O(\log\lambda)$-bit output length to break the original OWSGs. The output length is bounded by $O(\log\lambda)$ because the cardinality of $\gamma$-net in $m$-qubit space is bounded by $(C/\gamma)^{2^{2m}}$ for some constant $C$, which is polynomial for constant $\gamma$ and $m=\frac{1}{2}\log\log\lambda+O(1)$. For the OWSGs with pure state outputs, this can be improved to $\log\log\lambda+O(1).$
Finally, we study the regime of $\epsilon=1-1/{\mathrm{poly}}(\lambda)$ for some polynomial ${\mathrm{poly}}$. We refer to $(1-1/{\mathrm{poly}}(\lambda))$-OWSGs as weak OWSGs following [MY22a]. We show the following lower bound.
Theorem 1.4.
There do not exist weak OWSGs with $O(1)$-qubit outputs.
The proof is almost the same as that of Theorem 1.3 noting that $(C/\gamma)^{2^{2m}}$ is polynomial when $\gamma$ is inverse-polynomial and $m$ is constant.
Upper Bounds for ($\epsilon$)-OWSGs.
We show the above lower bounds are almost tight by showing almost matching upper bounds.
As already mentioned, the combination of existing works [BS20, CGG+23] gives $m(\lambda)$-qubit output OWSGs based on OWFs for any QPT-computable function $m(\lambda)=\omega(\log\lambda)$. This means that Theorem 3.1 is tight. We show that their construction can be extended to $\lambda^{-\Theta(1)}$-OWSGs to obtain the following theorem.
Theorem 1.5.
For any constant $c\in\mathbb{N}$, there exist $\lambda^{-c}$-OWSGs with $((c+1)\log\lambda+O(1))$-qubit pure state outputs assuming the existence of OWFs.
Comparing this with Theorem 1.2, we can see that it is almost tight though there is a small gap between $c\log\lambda-2$ and $(c+1)\log\lambda+O(1)$. The above theorem means that we can circumvent the barrier of $O(\log\lambda)$-qubit outputs for OWSGs (Theorem 1.1) if we slightly weaken (${\mathsf{negl}}(\lambda)$-)OWSGs to $\lambda^{-\Theta(1)}$-OWSGs.
Next, we show an upper bound for the output length of $\Theta(1)$-OWSGs.
Theorem 1.6.
For any constant $\epsilon>0$, there exist $\epsilon$-OWSGs with $O(\log\log\lambda)$-qubit pure state outputs assuming the existence of subexponentially secure OWFs.
Comparing this with Theorem 1.3, we can see that it is tight ignoring the constant factor for $\log\log\lambda$. We show this theorem by using quantum fingerprinting [BCWdW01]. Roughly, quantum finger printing encodes an $\ell$-bit classical string $x$ into an $O(\log(\ell)\log(\eta^{-1}))$-qubit pure state |$\ket{h_{x}}$ in such a way that $|\langle h_{x}|h_{x^{\prime}}\rangle|\leq\eta$ for any $x\neq x^{\prime}$. Our idea is to compose a OWF and quantum fingerprinting, i.e., our OWSG outputs |$\ket{h_{f(x)}}$ on input $x$ where $f$ is a OWF. When $\eta=\Theta(1)$, the length of the output is logarithmic in the output length of $f$. Thus, if we start from a OWF with ${\mathrm{polylog}}(\lambda)$-bit outputs, which exists assuming the existence of subexponentially secure OWFs, the output length of the above OWSG is $O(\log\log\lambda)$. We also show that the above construction preserves the one-wayness up to an additive constant error caused by quantum finger printing. This yields Theorem 1.6.
Finally, we show an upper bound for the output length of weak OWSGs.
Theorem 1.7.
For any QPT-computable function $m(\lambda)=\omega(1)$, there exist weak OWSGs with $m(\lambda)$-qubit pure state outputs assuming the existence of exponentially secure OWFs that map $\lambda$-bit inputs to $O(\lambda)$-bit outputs.
Comparing this with Theorem 1.4, we can see that this is tight. The main observation for proving this theorem is the following. For an angle $\theta$, let |$\ket{+_{\theta}}:=\frac{1}{\sqrt{2}}\left(\ket{0}+e^{i\theta}\ket{1}\right)$. Then we can show that for any distinct $y,y^{\prime}\in[\lambda]$, we have $|\langle+_{2\pi y/\lambda}|+_{2\pi y^{\prime}/\lambda}\rangle|\leq\sqrt{\frac{1+\cos(2\pi/\lambda)}{2}}=1-1/{\mathrm{poly}}(\lambda)$. This means that we can encode $\log\lambda$ classical bits into $1$ qubit if we allow the inner products to be as large as $1-1/{\mathrm{poly}}(\lambda)$. By repeating it, we can encode $t\log\lambda$ classical bits into $t$ qubits. By using this encoding instead of quantum fingerprinting in the construction for Theorem 1.6, we obtain weak OWSGs with $t$-qubit outputs starting from any OWFs with $t\log\lambda$-bit outputs. Assuming the existence of exponentially secure OWFs with linear expansion, there are OWFs with $t\log\lambda$-bit outputs for any super-constant $t$. This yields Theorem 1.7.
Our results on upper and lower bounds for ($\epsilon$-)OWSGs are summarized in Table 1.
Table 1: Summary of the results
Security of OWSGs Upperbound of output length Lowerbound of output length
[…]
(Assuming OWFs)
$\lambda^{-c}$ $(c+1)\log\lambda+O(1)$ [Theorem 1.5] $c\log\lambda-2$ [Theorem 1.2]
$(c=\Theta(1))$ (Assuming OWFs)
constant $O(\log\log\lambda)$ [Theorem 1.6] $\frac{1}{2}\log\log\lambda+O(1)$ [Theorem 1.3]
(Assuming subexp secure OWFs)
$1-\frac{1}{{\mathrm{poly}}(\lambda)}$ $\omega(1)$ [Theorem 1.7] $O(1)$[Theorem 1.4]
(Assuming exp secure OWFs)
EFIs.
We show matching lower and upper bounds for the length of EFIs. The lower bound is stated as follows.
Theorem 1.8.
There do not exist $O(\log\lambda)$-qubit EFIs.
We remark that the proof of this theorem is straightforward in the non-uniform setting: Since the two states that form an EFI are statistically far, there exists a measurement that distinguishes them with advantage $1/{\mathrm{poly}}(\lambda)$. If they are $O(\log\lambda)$-qubit states, the optimal distinguishing measurement can be described as a ${\mathrm{poly}}(\lambda)$-bit classical string. Thus, a non-uniform adversary that takes the description of the measurement as advice can break the computational indistinguishability. To extend it to the uniform setting, our idea is to use tomography. [^7] We thank Fermi Ma for suggesting the use of tomography. Roughly, we show that a uniform adversary can use tomography on the two states to efficiently approximate the optimal distinguishing measurement.
We show a matching upper bound as follows.
Theorem 1.9.
If exponentially secure PRGs exist, then $m(\lambda)$-qubit EFIs exist for any QPT computable function $m(\lambda)=\omega(\log\lambda)$.
The proof is almost obvious. By using an exponentially secure PRG $G:\{0,1\}^{m/2}\rightarrow\{0,1\}^{m}$, let $\rho_{0}:=\frac{1}{2^{m/2}}\sum_{x\in\{0,1\}^{m/2}}\ket{G(x)}\bra{G(x)}$ | and $\rho_{1}:=\frac{1}{2^{m}}\sum_{y\in\{0,1\}^{m}}\ket{y}\bra{y}$ |. Then $\rho_{0}$ and $\rho_{1}$ are computationally indistinguishable by exponential security of the PRG and $m(\lambda)=\omega(\log\lambda)$, but they are statistically far since the image size of $G$ is at most $2^{m/2}\ll 2^{m}$.
Quantum bit commitments.
It is known that EFIs exist if and only if quantum bit commitments exist [Yan22, BCQ23]. However, the conversion between them does not preserve the number of qubits, and thus the lower and upper bounds for the length of EFIs do not necessarily imply similar bounds for quantum bit commitments. To complement it, we also show similar lower and upper bounds for the commitment length (i.e., the number of qubits of the commitment register) of quantum bit commitments. That is, we show a $O(\log\lambda)$ lower bound and $\omega(\log\lambda)$ upper bound assuming the existence of exponentially secure PRGs. See Appendix A for the details.
1.2 Discussion
Our results suggest that the upper and lower bounds for the output length of ($\epsilon$-)OWSGs behave quite differently from those of OWFs.
  • It is easy to see that weak OWFs with $O(\log\lambda)$-bit outputs do not exist. On the other hand, assuming the existence of exponentially secure OWFs with linear expansion, we can show that there exist (strong) OWFs with $m$-bit outputs for any $m=\omega(\log\lambda)$ by a simple padding argument. This means that $O(\log\lambda)$ is the best lower bound for both OWFs and weak OWFs. On the other hand, our results suggest that the lower bounds for OWSGs and weak OWSGs are different. In particular, we show that there do not exist OWSGs with $O(\log\lambda)$-qubit outputs (Theorem 1.1), but there exist weak OWSGs with slightly super-constant-qubit outputs (Theorem 1.7), which circumvents the length barrier for standard OWSGs (Theorem 1.1).
  • For constructing OWFs with the optimal output length, i.e., those with slightly super-logarithmic-bit outputs, we need to assume exponential hardness of $\mathbf{NP}$ due to an obvious reason. On the other hand, we can construct OWSGs with slightly super-logarithmic-bit outputs only assuming polynomial hardness of OWFs [BS20, CGG+23]. Since we show that their construction is optimal in terms of output length (Theorem 1.1), this means that polynomial hardness is sufficient for achieving the optimal output length for OWSGs.
Our upper and lower bounds for the output length of OWSGs also imply those for several other primitives.
  1. 1.
    Quantum-public-key digital signatures: OWSGs are equivalent to (bounded poly-time secure) digital signatures with quantum public keys [MY22a]. Hence we have the following consequences. First, from Theorem 1.1, the length of the quantum public key of digital signatures cannot be $O(\log\lambda)$. Second, from Theorem 4.1, assuming OWFs, there exist one-time secure digital signatures with $m$-qubit quantum public keys for any $m=\omega(\log(\lambda))$. This can be regarded as a quantum advantage because it is unlikely that we can construct classical one-time secure digital signatures with similar public key length from (polynomially secure) OWFs.
  2. 2.
    Private-key quantum money: Private-key quantum money schemes with pure banknotes imply OWSGs [MY22a]. Hence, from Theorem 1.1, the length of pure quantum banknotes cannot be $O(\log\lambda)$.
Open problems.
Though we give matching upper and lower bounds for the output length of OWSGs and weak OWSGs, there are gaps for $\lambda^{\Theta(1)}$-OWSGs and $\Theta(1)$-OWSGs. For $\lambda^{c}$-OWSGs with $c=\Theta(1)$, our lower bound is $c\log\lambda-2$ (Theorem 1.2) whereas our upper bound is $(c+1)\log\lambda+O(1)$ (Theorem 1.5). For $\Theta(1)$-OWSGs, our lower bound is $(\log\log\lambda)/2+O(1)$ (Theorem 1.3) whereas upper bound is $O(\log\log\lambda)$ (Theorem 1.6). It is an interesting open problem to fill the gap.
Another interesting question is to construct quantum OWFs using short output OWSGs. Here, a quantum OWF is a quantum (pseudo)deterministic polynomial-time algorithm whose inputs and outputs are classical bit strings and the standard one-wayness is satisfied. The construction of quantum PRGs from short output PRSGs in [ALY23] would not be applied to the case of OWSGs, because both the proofs for (pseudodeterministic) correctness and security exploit (pseudo)randomness of output states of PRSGs. Answering this question provides a better understanding of quantum cryptographic primitives with short output lengths, and simplifies the proofs of Theorems 1.3 and 1.4.
1.3 Organization
We prepare definitions and technical tools in Section 2. We prove our results on lower bounds for the output length of ($\epsilon$-)OWSGs (Theorems 1.1, 1.2, 1.4 and 1.3) in Section 3. We prove our results on upper bounds for the output length of ($\epsilon$-)OWSGs (Theorems 1.6, 1.5 and 1.7) in Section 4. We show our results on lower and upper bounds for the length of EFIs (Theorems 1.8 and 1.9) in Section 5.
2 Preliminaries
Notations.
We use standard notations of quantum computing and cryptography. We use $\lambda$ as the security parameter. $[n]$ means the set $\{1,2,...,n\}$. For a finite set $S$, $x\leftarrow S$ means that an element $x$ is sampled uniformly at random from the set $S$. ${\mathsf{negl}}$ is a negligible function, and ${\mathrm{poly}}$ is a polynomial. PPT stands for (classical) probabilistic polynomial-time and QPT stands for quantum polynomial-time. We say that a function $m$ defined on $\mathbb{N}$ is QPT-computable if there is a QPT algorithm that computes $m(\lambda)$ on input $1^{\lambda}$ with probability $1$. We stress that the running time of the algorithm can be polynomial in $\lambda$ rather than in $\log\lambda$. For an algorithm $A$, $y\leftarrow A(x)$ means that the algorithm $A$ outputs $y$ on input $x$. For quantum states $\rho,\sigma$, $\|\rho-\sigma\|_{tr}$ means their trace distance. For any pure states |$\ket{\psi},\ket{\phi}$ it is known that $\||\psi\rangle\langle\psi|-|\phi\rangle\langle\phi|\|_{tr}=\sqrt{1-|\langle\psi|\phi\rangle|^{2}}$.
2.1 Basic Cryptographic Primitives
Definition 2.1 (OWFs).
A function $f:\{0,1\}^{*}\to\{0,1\}^{*}$ is a (quantum-secure) OWF if it is computable in classical deterministic polynomial-time, and for any QPT adversary $\mathcal{A}$,
$\Pr[f(x^{\prime})=f(x):x\leftarrow\{0,1\}^{\lambda},x^{\prime}\leftarrow\mathcal{A}(1^{\lambda},f(x))]\leq{\mathsf{negl}}(\lambda).$ (1)
We say that $f$ is a subexponentially secure OWF if there is a constant $10<c<$ 1 such that for any QPT adversary $\mathcal{A}$,
$\Pr[f(x^{\prime})=f(x):x\leftarrow\{0,1\}^{\lambda},x^{\prime}\leftarrow\mathcal{A}(1^{\lambda},f(x))]\leq 2^{-\lambda^{c}}$ (2)
for all sufficiently large $\lambda$.
We say that $f$ is an exponentially secure OWF if there is a constant $10<c<$ 1 such that for any QPT adversary $\mathcal{A}$,
$\Pr[f(x^{\prime})=f(x):x\leftarrow\{0,1\}^{\lambda},x^{\prime}\leftarrow\mathcal{A}(1^{\lambda},f(x))]\leq 2^{-c\lambda}$ (3)
for all sufficiently large $\lambda$.
Remark 2.2.
We require OWFs to be computable in classical deterministic polynomial-time following the standard definition. However, all of our positive results still follow even if we relax the requirement to only require them to be computable in QPT.
Definition 2.3 (PRGs).
A function $G:\{0,1\}^{*}\to\{0,1\}^{*}$ is a (quantum-secure) PRG if it is computable in classical deterministic polynomial-time, it maps $\lambda$-bit inputs to $\ell(\lambda)$-bit outputs for $\ell(\lambda)\geq\lambda+1$ and for any QPT adversary $\mathcal{A}$,
$\left|\Pr_{x\leftarrow\{0,1\}^{\lambda}}[1\leftarrow\mathcal{A}(1^{\lambda},G(x))]-\Pr_{y\leftarrow\{0,1\}^{\ell}}[1\leftarrow\mathcal{A}(1^{\lambda},y)]\right|\leq{\mathsf{negl}}(\lambda).$ (4)
We say that a PRG is exponentially secure if there exists a constant $10<c<$ 1 such that for any QPT adversary $\mathcal{A}$,
$\left|\Pr_{x\leftarrow\{0,1\}^{\lambda}}[1\leftarrow\mathcal{A}(1^{\lambda},G(x))]-\Pr_{y\leftarrow\{0,1\}^{\ell}}[1\leftarrow\mathcal{A}(1^{\lambda},y)]\right|\leq 2^{-c\lambda}.$ (5)
Definition 2.4 (One-way state generators (OWSGs)  [MY22a]).
For a function $\epsilon:\mathbb{N}\rightarrow[0,1]$, an $\epsilon$-one-way state generator ($\epsilon$-OWSG) is a tuple of QPT algorithms $(\mathsf{KeyGen},\mathsf{StateGen},\mathsf{Ver})$ that work as follows:
  • $\mathsf{KeyGen}(1^{\lambda})\to k:$ It is a QPT algorithm that, on input the security parameter $\lambda$, outputs a classical key $k$.
  • $\mathsf{StateGen}(k)\to\phi_{k}:$ It is a QPT algorithm that, on input $k$, outputs a quantum state $\phi_{k}$.
  • $\mathsf{Ver}(k^{\prime},\phi_{k})\to\top/\bot:$ It is a QPT algorithm that, on input $\phi_{k}$ and a bit string $k^{\prime}$, outputs $\top$ or $\bot$.
We require the following correctness and $\epsilon$-one-wayness.
Correctness:
$\Pr[\top\leftarrow\mathsf{Ver}(k,\phi_{k}):k\leftarrow\mathsf{KeyGen}(1^{\lambda}),\phi_{k}\leftarrow\mathsf{StateGen}(k)]\geq 1-{\mathsf{negl}}(\lambda).$
$\epsilon$-one-wayness:
[…]
$\Pr[\top\leftarrow\mathsf{Ver}(k^{\prime},\phi_{k}):k\leftarrow\mathsf{KeyGen}(1^{\lambda}),\phi_{k}\leftarrow\mathsf{StateGen}(k),k^{\prime}\leftarrow\mathcal{A}(1^{\lambda},\phi_{k}^{\otimes t})]\leq\epsilon(\lambda)$
for all sufficiently large $\lambda$.
We say that it is a OWSG if it is a $\lambda^{-c}$-OWSG for all $c\in\mathbb{N}$.
We say that it is a weak OWSG if there is $c\in\mathbb{N}$ such that it is a $(1-\lambda^{-c})$-OWSG.
We say that it has $m(\lambda)$-qubit outputs if $\phi_{k}$ is an $m(\lambda)$-qubit state for all $k$ in the support of $\mathsf{KeyGen}(1^{\lambda})$.
We say that it has pure output states if $\phi_{k}$ is a pure state for all $k$ in the support of $\mathsf{KeyGen}(1^{\lambda})$.
Definition 2.5 (Pseudo-random quantum states generators (PRSGs) [JLS18]).
For a function $\epsilon:\mathbb{N}\to[0,1]$, An $\epsilon$-pseudorandom quantum states generator ($\epsilon$-PRSG) is a tuple of QPT algorithms $(\mathsf{KeyGen},\mathsf{StateGen})$ that work as follows:
  • $\mathsf{KeyGen}(1^{\lambda})\to k:$ It is a QPT algorithm that, on input the security parameter $\lambda$, outputs a classical key $k$.
  • $\mathsf{StateGen}(k)\to\ket{\phi_{k}}:$ It is a QPT algorithm that, on input $k$, outputs an $m$-qubit quantum state |$\ket{\phi_{k}}$.
We require the following pseudorandomness: For any polynomial $t$ and any QPT algorithm $\mathcal{A}$,
$\left|\Pr_{k\leftarrow\mathsf{KeyGen}(1^{\lambda})}\left[1\leftarrow\mathcal{A}\left(1^{\lambda},\ket{\phi_{k}}^{\otimes t}\right)\right]-\Pr_{\ket{\psi}\leftarrow\mu_{m}}\left[1\leftarrow\mathcal{A}\left(1^{\lambda},\ket{\psi}^{\otimes t}\right)\right]\right|\leq\epsilon(\lambda)$
for all sufficiently large $\lambda$, where $\mu_{m}$ is the Haar measure on $m$-qubit quantum states. We say that it is a PRSG if it is a $\lambda^{-c}$-PRSG for all $c\in\mathbb{N}$.
Definition 2.6 (EFIs [BCQ23]).
An EFI is an algorithm $\mathsf{StateGen}(b,1^{\lambda})\to\rho_{b}$ that, on input $b\in\{0,1\}$ and the security parameter $\lambda$, outputs a quantum state $\rho_{b}$ such that all of the following three conditions are satisfied.
  • It is a uniform QPT algorithm.
  • $\rho_{0}$ and $\rho_{1}$ are computationally indistinguishable. In other words, for any QPT adversary $\mathcal{A}$, $|\Pr[1\leftarrow\mathcal{A}(1^{\lambda},\rho_{0})]-\Pr[1\leftarrow\mathcal{A}(1^{\lambda},\rho_{1})]|\leq{\mathsf{negl}}(\lambda).$
  • $\rho_{0}$ and $\rho_{1}$ are statistically distinguishable, i.e., $\|\rho_{0}-\rho_{1}\|_{tr}\geq\frac{1}{{\mathrm{poly}}(\lambda)}.$
We call it an $\ell(\lambda)$-qubit EFI if $\rho_{0}$ and $\rho_{1}$ are $\ell(\lambda)$-qubit states.
Remark 2.7.
In the original definition of EFIs in [BCQ23], computational indistinguishability is required to hold against non-uniform QPT adversaries whereas we only require computational indistinguishability against uniform QPT adversaries. This only makes our result on the lower bound for the output length of EFIs stronger.
2.2 Useful Lemmata
Lemma 2.8 (Welch Bound [Wel74, BS08]).
Let $n,d,k$ be positive integers. For any family {|$\{\ket{\psi_{i}}\}_{i\in[n]}$ of (not necessarily unit) vectors in $\mathbb{C}^{d}$, it holds that
(${d+k-1\choose k}\sum_{i,j\in[n]}|\langle\psi_{i}|\psi_{j}\rangle|^{2k}\geq\left(\sum_{i\in[n]}\langle\psi_{i}|\psi_{i}\rangle^{k}\right)^{2}.$
In particular, we use the following corollary.
Corollary 2.9.
Let $n,m$ be positive integers. For any family {|$\{\ket{\phi_{i}}\}_{i\in[n]}$ of (normalized) $m$-qubit pure states and any distribution $\mathcal{D}$ over $[n]$, it holds that
$\mathbb{E}_{i\leftarrow\mathcal{D},j\leftarrow\mathcal{D}}|\langle\phi_{i}|\phi_{j}\rangle|^{2}\geq 2^{-m}.$
Proof.
For each $i\in[n]$, let $p_{i}:=\Pr[i\leftarrow\mathcal{D}]$ and |$\ket{\psi_{i}}:=\sqrt{p_{i}}\ket{\phi_{i}}$. Applying Lemma 2.8 for {|$\{\ket{\psi_{i}}\}_{i\in[n]}$ where $k:=1$ and $d:=2^{m}$, we obtain
$2^{m}\sum_{i,j\in[n]}p_{i}p_{j}|\langle\phi_{i}|\phi_{j}\rangle|^{2}\geq\left(\sum_{i\in[n]}p_{i}\langle\phi_{i}|\phi_{i}\rangle\right)^{2}.$
By definition, we have
$\sum_{i,j\in[n]}p_{i}p_{j}|\langle\phi_{i}|\phi_{j}\rangle|^{2}=\mathbb{E}_{i\leftarrow\mathcal{D},j\leftarrow\mathcal{D}}|\langle\phi_{i}|\phi_{j}\rangle|^{2}.$
On the other hand, since |$\ket{\phi_{i}}$ is a normalized state, we have
$\left(\sum_{i\in[n]}p_{i}\langle\phi_{i}|\phi_{i}\rangle\right)^{2}=\left(\sum_{i\in[n]}p_{i}\right)^{2}=1^{2}=1.$
Combining the above, we obtain Corollary 2.9. ∎
Lemma 2.10 (Quantum state tomography [ALY23, Corollary 2.16]).
Let $\delta=\delta(\lambda)\in(0,1]$ and $d=d(\lambda)\in\mathbb{N}$. There exists a tomography procedure $\mathsf{Tomography}$ that takes $t(\lambda):=144\lambda d^{4}/\delta^{2}$ copies of a $d$-dimensional mixed quantum state $\rho$, and outputs a Hermitian matrix $M\in\mathbb{C}^{d\times d}$ such that
$\Pr\left[\|M-\rho\|_{tr}\leq\delta:M\leftarrow\mathsf{Tomography}(\rho^{\otimes t},\delta)\right]\geq 1-{\mathsf{negl}}(\lambda).$
Moreover, the running time of $\mathsf{Tomography}$ is polynomial in $1/\delta,d$ and $\lambda$.
Remark 2.11.
[ALY23, Corollary 2.16] does not mention that the output matrix is Hermitian. This fact can be checked by following the proof; the starting point is the folklore tomography (see [Low21, Section 1.5.3]) that approximates $\rho=\frac{1}{d}\sum\mathrm{Tr}(P_{i}\rho)P_{i}$ for Pauli matrices $P_{i}$ by $N=\sum\mu_{i}P_{i}$ for $\mu_{i}\approx\frac{1}{d}\mathrm{Tr}(P_{i}\rho)$. The matrix $N$ is Hermitian, and the above lemma, which is a trace-distance version of [AGQY22, Corollary 7.6], only chooses a nice $M$ among many independent $N$’s, so that $M$ is still Hermitian.
Lemma 2.12 ($\gamma$-net of quantum states [AS17, Section 5]).
Let $\gamma\in(0,1]$ and $d\in\mathbb{N}$. There exist some constant $C>0$ and a set $\mathsf{Net}(d,\gamma)$ of density matrices, called an $\gamma$-net (of quantum states), of size $|\mathsf{Net}(d,\gamma)|\leq(C/\gamma)^{d^{2}}$ such that the following holds: For any $d$-dimensional mixed quantum state $\rho$, there exists $N\in\mathsf{Net}(d,\gamma)$ such that $\|N-\rho\|_{tr}\leq\gamma$. Moreover, there exists an explicit construction of $\mathsf{Net}(d,\gamma)$ in time $|\mathsf{Net}(d,\gamma)|\cdot{\mathrm{poly}}(d,1/\gamma)$. If we only consider the pure states, $|\mathsf{Net}(d,\gamma)|\leq(C/\gamma)^{d}$ holds.
Sketch of proof.
By purification, it suffices to show that there exists an explicit $\gamma$-net for pure states in dimension $d^{2}$ of size $(C/\gamma)^{d^{2}}$. This is given in [AS17, Exercise 5.22] by normalizing an (explicit) $\gamma$-net in the Boolean cube for Hamming distance in the same dimension. ∎
Lemma 2.13 (Quantum Fingerprinting [BCWdW01]).
There is a ${\mathrm{poly}}(\ell)\log\eta^{-1}$-time quantum algorithm that takes $x\in\{0,1\}^{\ell}$ as input and outputs an $O(\log(\ell)\log(\eta^{-1}))$-qubit state |$\ket{h_{x}}$ such that $|\langle h_{x}|h_{x^{\prime}}\rangle|\leq\eta$ for each $x\neq x^{\prime}$.
Proof.
Because the above statement is not explicitly written in [BCWdW01], we here provide its proof. Let $E:\{0,1\}^{\ell}\to\{0,1\}^{m}$ be an error-correcting code such that $m=c\ell$ and the Hamming distance between any two distinct code words $E(x)$ and $E(y)$ (i.e., the number of bit positions where they differ) is at least $(1-\delta)m$, where $c$ and $\delta$ are positive constants. For some specific codes, we can choose any $c>2$ and have $\delta<\frac{9}{10}+\frac{1}{15c}$ for sufficiently large $\ell$ [BCWdW01]. Let us take $c=4$. Then $\delta<\frac{9}{10}+\frac{1}{15c}=\frac{11}{12}$ for sufficiently large $\ell$. Let $r\coloneqq\lfloor\frac{\log\eta}{\log\frac{11}{12}}\rfloor+1$. For each $x\in\{0,1\}^{\ell}$, define the state
|$\ket{h_{x}}\coloneqq\left(\frac{1}{\sqrt{m}}\sum_{i=1}^{m}\ket{i}\ket{E_{i}(x)}\right)^{\otimes r},$ (6)
where $E_{i}(x)$ is the $i$th bit of $E(x)$. Then
$|\langle h_{x}|h_{x^{\prime}}\rangle|=\left(\frac{1}{m}\sum_{i=1}^{m}\langle E_{i}(x)|E_{i}(x^{\prime})\rangle\right)^{r}\leq\delta^{r}<\left(\frac{11}{12}\right)^{\frac{\log\eta}{\log\frac{11}{12}}}=\eta.$ (7)
The number $N$ of qubits of |$\ket{h_{x}}$ is
$N=r(\lceil\log(m)\rceil+1)\leq\left(\frac{\log\eta^{-1}}{\log\frac{12}{11}}+1\right)(\log(c\ell)+2)<(8\log\eta^{-1}+1)(\log\ell+4).$ (8)
Therefore we have $N=O(\log\ell\log\eta^{-1})$.
Remark 2.14.
The work [BCWdW01, Section 3] also shows an alternative non-constructive method for quantum fingerprinting with better output length. With some rephrasing, they show that we can take $m=\log\ell+O(\log\eta^{-1})$ in such a way that if we define |$\ket{h_{x}}:=2^{-m/2}\sum_{z\in\{0,1\}^{m}}(-1)^{H(x||z)}\ket{z}$ for a random function $H:\{0,1\}^{\ell+m}\rightarrow\{0,1\}$, then for $(1-2^{-\Omega(\ell)})$-fraction of $H$, we have $|\langle h_{x}|h_{x^{\prime}}\rangle|\leq\eta$ for all pairs of distinct $\ell$-bit strings $x\neq x^{\prime}$. Then they argue that there exists a fixed $H$ which gives a quantum fingerprinting with the desired parameter by a non-constructive argument. We observe that the above construction can be directly used as quantum fingerprinting in the quantum random oracle model (QROM) [BDF+11]. This means that we have a quantum fingerprinting with output length $\log\ell+O(\log\eta^{-1})$ in the QROM, which is better than the one given in Lemma 2.13. We do not know how to achieve the same bound in the plain model (i.e., without relying on random oracles). The improvement from $O(\log(\ell)\log(\eta^{-1}))$ to $\log\ell+O(\log\eta^{-1})$ does not lead to any asymptotical improvement for output length of $\epsilon$-OWSGs in the parameter regime which we consider in this paper.
Lemma 2.15 ([Ras07, Lemma 4]).
For any Hermitian matrices $A$ and $B$, it holds that
$\max_{\Pi}\mathrm{Tr}(\Pi(A-B))=\|A-B\|_{tr}+\frac{\mathrm{Tr}(A-B)}{2}$
where the maximum is taken over all projections $\Pi$.
Corollary 2.16.
For any Hermitian matrices $A$ and $B$ and any projection $\Pi$, it holds that
$\left|\mathrm{Tr}\left(\Pi(A-B)\right)\right|\leq 2\|A-B\|_{tr}.$
Proof.
It immediately follows from Lemma 2.15 and $\mathrm{Tr}(M)\leq\|M\|_{1}=2\|M\|_{tr}$ for any Hermitian matrix $M$. ∎
Lemma 2.17.
Let {|$\{\ket{\psi_{k}}\}_{k\in\{0,1\}^{n}}$ be any set of pure states. Then
$F\left(\frac{1}{2^{n}}\sum_{k\in\{0,1\}^{n}}\ket{\psi_{k}}\bra{\psi_{k}},\frac{I^{\otimes m}}{2^{m}}\right)\leq\frac{2^{n}}{2^{m}}.$ (9)
Proof.
Let $\sum_{k\in\{0,1\}^{n}}\alpha_{k}\ket{\phi_{k}}\bra{\phi_{k}}$ | be a diagonalization of $\frac{1}{2^{n}}\sum_{k\in\{0,1\}^{n}}\ket{\psi_{k}}\bra{\psi_{k}}$ |. Then
$F\left(\frac{1}{2^{n}}\sum_{k\in\{0,1\}^{n}}\ket{\psi_{k}}\bra{\psi_{k}},\frac{I^{\otimes m}}{2^{m}}\right)=\left(\sum_{k}\sqrt{\frac{\alpha_{k}}{2^{m}}}\right)^{2}\leq(\sum_{k}\alpha_{k})\left(\sum_{k}\frac{1}{2^{m}}\right)=\frac{2^{n}}{2^{m}}.$ (10)
Here in the inequality we have used Cauchy–Schwarz inequality. ∎
3 Lower Bounds for ($\epsilon$-)OWSGs
In this section, we prove lower bounds for output length of ($\epsilon$-)OWSGs (Theorems 1.1, 1.2, 1.4 and 1.3).
3.1 Lower Bounds for OWSGs and $\lambda^{-\Theta(1)}$-OWSGs
Theorem 3.1 (Restatement of Theorem 1.1).
There do not exist OWSGs with $O(\log\lambda)$-qubit outputs.
Proof.
Suppose that QPT algorithms $(\mathsf{KeyGen},\mathsf{StateGen},\mathsf{Ver})$ satisfy the syntax and correctness of OWSGs and has $m=O(\log\lambda)$-qubit outputs. We first prove the following:
$\mathbb{E}_{k,k^{\prime}}\|\phi_{k}-\phi_{k^{\prime}}\|_{tr}\leq\sqrt{1-2^{-m}}.$ (11)
We prove this below. For each $k$, we consider the diagonalization of $\phi_{k}$ to write
$\phi_{k}=\sum_{i}p_{k,i}|\phi_{k,i}\rangle\langle\phi_{k,i}|.$
Then we have
$\mathbb{E}_{k,k^{\prime}}\|\phi_{k}-\phi_{k^{\prime}}\|_{tr}$ $=\mathbb{E}_{k,k^{\prime}}\left\|\sum_{i}p_{k,i}|\phi_{k,i}\rangle\langle\phi_{k,i}|-\sum_{j}p_{k^{\prime},j}|\phi_{k^{\prime},j}\rangle\langle\phi_{k^{\prime},j}|\right\|_{tr}$ (12)
$=\mathbb{E}_{k,k^{\prime}}\left\|\sum_{i,j}p_{k,i}p_{k^{\prime},j}\left(|\phi_{k,i}\rangle\langle\phi_{k,i}|-|\phi_{k^{\prime},j}\rangle\langle\phi_{k^{\prime},j}|\right)\right\|_{tr}$ (13)
$\leq\mathbb{E}_{k,k^{\prime}}\sum_{i,j}p_{k,i}p_{k^{\prime},j}\left\||\phi_{k,i}\rangle\langle\phi_{k,i}|-|\phi_{k^{\prime},j}\rangle\langle\phi_{k^{\prime},j}|\right\|_{tr}$ (14)
$=\mathbb{E}_{k,k^{\prime}}\sum_{i,j}p_{k,i}p_{k^{\prime},j}\sqrt{1-|\langle\phi_{k,i}|\phi_{k^{\prime},j}\rangle|^{2}}$ (15)
$\leq\sqrt{1-\mathbb{E}_{k,k^{\prime}}\sum_{i,j}p_{k,i}p_{k^{\prime},j}|\langle\phi_{k,i}|\phi_{k^{\prime},j}\rangle|^{2}}$ (16)
$\leq\sqrt{1-2^{-m}},$ (17)
where Equation 14 follows from the triangle inequality for trace distance, Equation 16 follows from the concavity of the function $\sqrt{1-x}$, and Equation 17 follows from Corollary 2.9. This completes the proof of Equation 11.
Then we consider a trivial adversary $\mathcal{A}$ against its one-wayness that ignores its input (except for the security parameter) and simply outputs a fresh random key $k^{\prime}\leftarrow\mathsf{KeyGen}(1^{\lambda})$. Then we have
$\Pr[\mathcal{A}~{}\text{wins}]$ $=\mathbb{E}_{k,k^{\prime}}\Pr[\mathsf{Ver}(k^{\prime},\phi_{k})=\top]$ (18)
$\geq\mathbb{E}_{k,k^{\prime}}\left[\Pr[\mathsf{Ver}(k,\phi_{k})=\top]-\|\phi_{k}-\phi_{k^{\prime}}\|_{tr}\right]$ (19)
$\geq 1-{\mathsf{negl}}(\lambda)-\sqrt{1-2^{-m}},$ (20)
where Equation 20 follows from the correctness of $(\mathsf{KeyGen},\mathsf{StateGen},\mathsf{Ver})$ and Equation 11. Note that we have $\sqrt{1-x}\leq 1-x/2$ for all $0\leq x\leq 1$. Therefore we have
$\Pr[\mathcal{A}~{}\text{wins}]\geq 2^{-m-1}-{\mathsf{negl}}(\lambda).$ (21)
Since $m=O(\log\lambda)$, the RHS is non-negligible and thus $\mathcal{A}$ succeeds in breaking one-wayness. ∎
Remark 3.2.
In the above proof, $\mathcal{A}$ does not use the target state at all. This means that even zero-copy OWSGs (where the adversary is given no copy of the target state) with $O(\log\lambda)$-qubit outputs do not exist.
The above proof can be easily extended to give a lower bound for $\lambda^{-\Theta(1)}$-OWSGs.
Theorem 3.3 (Restatement of Theorem 1.2).
For any constant $c\in\mathbb{N}$, there do not exist $\lambda^{-c}$-OWSGs of which output length is at most $c\log\lambda-2$.
Proof.
Suppose that the output length is $m<c^{\prime}\log\lambda-2$. By the proof of Theorem 3.1, there is a QPT adversary that wins with probability at least $2^{-m-1}-{\mathsf{negl}}(\lambda)\geq 2\lambda^{-c}-{\mathsf{negl}}(\lambda)\geq\lambda^{-c}$ for all sufficiently large $\lambda$. This means that it is not a $\lambda^{-c}$-OWSG. ∎
Remark 3.4.
As one can easily see from the proof, we can improve the lower bound from $c\log\lambda-2$ to $c\log\lambda-1-\epsilon$ for any constant $\epsilon>0$.
3.2 Lower Bounds for $\Theta(1)$-OWSGs and Weak OWSGs.
For proving Theorems 1.4 and 1.3, we prove the following theorem.
Theorem 3.5.
Let $\Delta=\Delta(\lambda)>0$ and $m=m(\lambda)\in\mathbb{N}$ and $C$ be a constant taken from Lemma 2.12. If $(6C/\Delta)^{2^{2m}}={\mathrm{poly}}(\lambda)$, there do not exist $(1-\Delta)$-OWSGs with $m$-qubit outputs.
Proof.
Suppose that QPT algorithms $(\mathsf{KeyGen},\mathsf{StateGen},\mathsf{Ver})$ satisfy the syntax and correctness of $(1-\Delta)$-OWSGs with $m$-qubit outputs. Let $\gamma=\Delta/6$, $d=2^{m}$, $t=144\lambda d^{4}/\gamma^{2}$, and $T=\frac{(C/\gamma)^{d^{2}}\log\lambda}{\gamma}$. We consider the following adversary $\mathcal{A}(1^{\lambda},\phi_{k}^{\otimes t})$:
  1. 1.
    Construct an $\gamma$-net $\mathsf{Net}(d,\gamma)$.
  2. 2.
    Given $\phi_{k}^{\otimes t}$, run $M\leftarrow\mathsf{Tomography}(\phi_{k}^{\otimes t},\gamma)$.
  3. 3.
    Construct a set $\mathcal{N}=\{N\in\mathsf{Net}(d,\gamma):\|N-M\|_{tr}\leq\gamma\}.$
  4. 4.
    For each $i\in[T]$, do:
    1. (a)
      Sample $k_{i}\leftarrow\mathsf{KeyGen}(1^{\lambda})$ and run $M_{i}\leftarrow\mathsf{Tomography}(\phi_{k_{i}}^{\otimes t},\gamma).$
    2. (b)
      If there exists $N\in\mathcal{N}$ such that $\|M_{i}-N\|_{tr}\leq\gamma$, output $k^{\prime}:=k_{i}$ and terminate.
  5. 5.
    Output $\bot$.
Because of Lemmata 2.10 and 2.12, the time complexity of this algorithm is bounded by a polynomial in $1/\gamma,d,\lambda,T$ and $|\mathsf{Net}(d,\gamma)$. By the condition, $|\mathsf{Net}(d,\gamma)|\leq(C/\gamma)^{d^{2}}=(6C/\Delta)^{2^{2m}}={\mathrm{poly}}(\lambda)$ and this implies that $1/\Delta={\mathrm{poly}}(\lambda)$ and $d=2^{m}={\mathrm{poly}}(\lambda)$. This proves that $\mathcal{A}$ is a QPT algorithm.
To prove the correctness, we first argue that $\mathcal{A}$ outputs $\bot$ with a small probability. Let $\epsilon_{bad}:=\frac{\gamma}{|\mathsf{Net}(d,\gamma)|}$. Let $\mathcal{B}$ be the set of elements $N\in\mathsf{Net}(d,\gamma)$ such that the following inequality holds:
$\Pr\left[\|N-M^{\prime}\|_{tr}\leq\gamma:k^{\prime}\leftarrow\mathsf{KeyGen}(1^{\lambda}),M^{\prime}\leftarrow\mathsf{Tomography}(\phi_{k^{\prime}}^{\otimes t},\gamma)\right]\leq\epsilon_{bad}.$ (22)
Let $\sf{bad}$ be the event that $\mathcal{N}\cap\mathcal{B}\neq\emptyset$. It holds that
$\Pr[\sf{bad}]=\Pr[\mathcal{N}\cap\mathcal{B}\neq\emptyset]\leq\sum_{N\in\mathcal{B}}\Pr[N\in\mathcal{N}]\leq|\mathcal{B}|\cdot\epsilon_{bad}\leq\gamma,$ (23)
where we have used $|\mathcal{B}|\leq|\mathsf{Net}(d,\gamma)|$ at the last inequality.
Suppose that $\sf{bad}$ does not occur. Then, since $\mathcal{N}$ is not empty, there exists an element $N\in\mathcal{N}$ such that
$\Pr\left[\|N-M^{\prime}\|_{tr}\leq\gamma:k^{\prime}\leftarrow\mathsf{KeyGen}(1^{\lambda}),M^{\prime}\leftarrow\mathsf{Tomography}(\phi_{k^{\prime}}^{\otimes t},\gamma)\right]>\epsilon_{bad},$ (24)
which means that each iteration of the fourth step terminates with probability at least $\frac{\gamma}{|\mathsf{Net}(d,\gamma)|}$. Then we have
$\Pr\left[k^{\prime}\neq\bot:k\leftarrow\mathsf{KeyGen}(1^{\lambda}),k^{\prime}\leftarrow\mathcal{A}(1^{\lambda},\phi_{k}^{\otimes t})\right]$ (25)
$\geq\Pr\left[k^{\prime}\neq\bot:k\leftarrow\mathsf{KeyGen}(1^{\lambda}),k^{\prime}\leftarrow\mathcal{A}(1^{\lambda},\phi_{k}^{\otimes t})|\lnot{\sf{bad}}\right]\cdot\Pr[\lnot{\sf{bad}}]$ (26)
$\geq\left(1-\left(1-\epsilon_{bad}\right)^{T}\right)\cdot(1-\gamma)\geq 1-\gamma-{\mathsf{negl}}(\lambda),$ (27)
where the first inequality in Equation 27 holds because the iterations are independent, and the last inequality holds because of the choice of $T\geq{\log\lambda}/{\epsilon_{bad}}$.
Finally, we prove that the output of $\mathcal{A}$ is accepted by $\mathsf{Ver}$ with high probability whenever it is not $\bot$. If $\mathcal{A}$ outputs $k_{i}$, $\|\phi_{k}-\phi_{k_{i}}\|_{tr}\leq 4\gamma$ with probability $1-{\mathsf{negl}}(\lambda)$. By the triangular inequality, we have
$\|\phi_{k}-\phi_{k_{i}}\|_{tr}\leq\|\phi_{k}-M\|_{tr}+\|M-N\|_{tr}+\|N-M_{i}\|_{tr}+\|M_{i}-\phi_{k_{i}}\|_{tr},$ (28)
all of which are bounded by $\gamma$ with probability $1-{\mathsf{negl}}(\lambda)$ because of Lemma 2.10 and the definition of $\mathsf{Net}(d,\gamma)$. This implies that
$\Pr[\top\leftarrow\mathsf{Ver}(k_{i},\phi_{k})]\geq\Pr[\top\leftarrow\mathsf{Ver}(k,\phi_{k})]-\|\phi_{k}-\phi_{k_{i}}\|_{tr}\geq 1-4\gamma-{\mathsf{negl}}(\lambda)$ (29)
holds with probability $1-{\mathsf{negl}}(\lambda)$, where the last inequality follows from the correctness of $(\mathsf{KeyGen},\mathsf{StateGen},\mathsf{Ver})$. Using $\gamma=\Delta/6$, the advantage of $\mathcal{A}$ is bounded as follows:
[…]
$\geq(1-4\gamma-{\mathsf{negl}}(\lambda))\cdot(1-\gamma-{\mathsf{negl}}(\lambda))\cdot(1-{\mathsf{negl}}(\lambda)$ (31)
$\geq 1-5\gamma-{\mathsf{negl}}(\lambda)\geq 1-\Delta$ (32)
where the first term of Equation 31 is from Equation 29, the second term is from Equation 27, and the last term is the probability that Equation 29 holds. Therefore, $\mathcal{A}$ is a QPT algorithm breaking $(1-\Delta)$-OWSG of $(\mathsf{KeyGen},\mathsf{StateGen},\mathsf{Ver})$. ∎
Theorems 1.4 and 1.3 follow from Theorem 3.5 as immediate corollaries.
Corollary 3.6 (Restatement of Theorem 1.3).
There do not exist $\epsilon$-OWSGs with $m$-qubit outputs for any constant $\epsilon>0$ and $m=\frac{1}{2}\log\log\lambda+O(1)$.
Proof.
If the output length $m$ is $m=\frac{1}{2}\log\log\lambda+O(1)$ so that $2^{2m}=O(\log\lambda),$ then for any constant $\epsilon>0$, $(6C/(1-\epsilon))^{2^{2m}}=(6C/(1-\epsilon))^{O(\log\lambda)}={\mathrm{poly}}(\lambda)$ and we can apply Theorem 3.5. ∎
Remark 3.7.
For the pure case, we can prove that there do not exist $\epsilon$-OWSGs with $(\log\log\lambda+O(1))$-qubit outputs for any constant $\epsilon>0$ using the pure case of Lemma 2.12.
Corollary 3.8 (Restatement of Theorem 1.4).
There do not exist weak OWSGs with any constant output length.
Proof.
Consider the security of weak OWSGs with the constant output length $m$, which is $(1-\epsilon)$-secure for some inverse polynomial $\epsilon$. Since $m$ is constant, $(6C/\epsilon)^{2^{2m}}={\mathrm{poly}}(\lambda)$, thus Theorem 3.5 applies. ∎
4 Upper Bounds for ($\epsilon$-)OWSGs
In this section, we prove upper bounds for output length of ($\epsilon$-)OWSGs (Theorems 1.5, 1.6 and 1.7).
4.1 Upper Bounds for OWSGs and $\lambda^{-\Theta(1)}$-OWSGs
We first observe that the combination of previous results already gives OWSGs with slightly super-logarithmic-qubit outputs.
Theorem 4.1 ([BS20, CGG+23]).
If there exist OWFs, then there exist OWSGs with $m(\lambda)$-qubit pure state outputs for any QPT-computable function $m(\lambda)=\omega(\log\lambda)$.
Proof.
According to [BS20], if there exist OWFs, then there exist PRSGs with input length $\lambda$ and any output length $m$. From [CGG+23, Theorem 3.4], these PRSGs with $m=\omega(\log\lambda)$ are also OWSGs. ∎
We show that the result of [CGG+23] can be adapted to construct $1/{\mathrm{poly}}(\lambda)$-OWSGs with $O(\log\lambda)$-qubit outputs. We recall the following lemma from [CGG+23, Lemma 3.2] with a small modification.
Lemma 4.2.
Let $(\mathsf{KeyGen},\mathsf{StateGen})$ be an $\epsilon$-PRSG, where $\mathsf{StateGen}$ maps $n$-bit strings to $m$-qubit pure states. Let $h\in[0,1]$ and $\delta=2^{n}\cdot\left(1-h\right)^{2^{m}-1}+h$. Define a QPT algorithm $\mathsf{Ver}(k^{\prime},\ket{\phi_{k}})\to\top/\bot$ to be the projection of |$\ket{\phi_{k}}$ onto |$\ket{\phi_{k^{\prime}}}$ with the output $\top$ only when the projection is successful. Then $(\mathsf{KeyGen},\mathsf{StateGen},\mathsf{Ver})$ is an $(\epsilon+\delta)$-OWSG.
Proof sketch of Lemma 4.2.
Most parts of the proof are identical to the original one. When proving the security, we recall the following concentration inequality for Haar random states instead of [CGG+23, Lemma 3.1].
Lemma 4.3.
For any $h\in[0,1]$ and for any $m$-qubit quantum state $\phi_{0}$, it holds that for the Haar measure on $m$-qubit quantum states $\mu_{m}$,
$\Pr_{\ket{\psi}\leftarrow\mu_{m}}\left[|\braket{\psi}{\phi_{0}}|^{2}\geq h\right]=(1-h)^{2^{m}-1}.$
The proof of Lemma 4.3 can be found in the proof of [AK07, Lemma 3.6]. Using Lemma 4.3 instead of [CGG+23, Lemma 3.1] gives the choice $\delta=2^{n}\cdot(1-h)^{2^{m}-1}+h$, where $h$ has a similar role to $1/f(n)$ in the original proof. ∎
Theorem 4.4 (Restatement of Theorem 1.5).
For any constant $c\in\mathbb{N}$, there exist $\lambda^{-c}$-OWSGs with $(c+1)\log\lambda+O(1)$-qubit pure state outputs assuming the existence of OWFs.
Proof.
By [BS20], there exist PRSGs with input length $\lambda$ and any output length $m$ assuming the existence of OWFs. We apply Lemma 4.2 with $h=\lambda^{-c}/4$ and
$m\geq\log\left(\frac{\log_{e}\left(2^{\lambda+2}\lambda^{c}\right)}{h}+1\right)~{}\Longrightarrow~{}2^{n}\cdot(1-h)^{2^{m}-1}\leq\lambda^{-c}/4,$ (33)
which gives $\delta=\lambda^{-c}/2$. To prove the asymptotic bound of $m(\lambda)$, we choose a constant $K$ such that $2^{\lambda+2}\lambda^{c}\leq e^{\lambda}$ for all $\lambda\geq K$. In this case, the minimal $m$ satisfies
$m\geq\log\left(4\lambda^{c+1}+1\right)\geq(c+1)\log\lambda+3,$ (34)
thus we can choose $m(\lambda)=(c+1)\log\lambda+O(1)$. ∎
4.2 Upper Bound for $\Theta(1)$-OWSGs
For proving Theorem 1.6, we prove the following theorem.
Theorem 4.5.
If there exist OWFs that map $\lambda$-bit inputs to $\ell(\lambda)$-bit outputs, then there exist $(\eta(\lambda)^{2}+{\mathsf{negl}}(\lambda))$-OWSGs with $O(\log(\ell(\lambda))\log(\eta(\lambda)^{-1}))$-qubit pure state outputs for any QPT-computable function $\eta(\lambda)$.
Proof.
For notational simplicity, we simply write $\ell$ and $\eta$ to mean $\ell(\lambda)$ and $\eta(\lambda)$, respecitively. Let $f$ be a OWF that maps $\lambda$-bit inputs to $\ell$-bit outputs. We construct an $(\eta^{2}+{\mathsf{negl}}(\lambda))$-OWSG as follows:
  • $\mathsf{KeyGen}(1^{\lambda})\to k:$ Sample $k\leftarrow\{0,1\}^{\lambda}$.
  • $\mathsf{StateGen}(k)\to\ket{\phi_{k}}:$ Output |$\ket{\phi_{k}}\coloneqq\ket{h_{f(k)}}$, where |$\ket{h_{f(k)}}$ is the quantum fingerprinting state of $f(k)$ given in Lemma 2.13.
  • $\mathsf{Ver}(k^{\prime},\ket{\phi_{k}})\to\top/\bot:$ Project |$\ket{\phi_{k}}$ onto |$\ket{\phi_{k^{\prime}}}$. If the projection is successful, output $\top$. Otherwise, output $\bot$.
Let $\mathcal{A}$ be a QPT adversary of this OWSG that receives $t$ copies of |$\ket{\phi_{k}}$. Then, as we will show below, there exists a QPT adversary $\mathcal{B}$ of the OWF $f$ that satisfies [^9] We say that $\mathcal{A}$ wins if $\mathsf{Ver}(k^{\prime},\ket{\phi_{k}})\rightarrow\top$ where $k\leftarrow\mathsf{KeyGen}(1^{\lambda})$ and $k^{\prime}\leftarrow\mathcal{A}(1^{\lambda},\ket{\phi_{k}}^{\otimes t})$. Similarly, we say that $\mathcal{B}$ wins if $f(x^{\prime})=f(x)$ where $x\leftarrow\{0,1\}^{\lambda}$ and $x^{\prime}\leftarrow\mathcal{B}(1^{\lambda},f(x))$.
$\Pr[\mathcal{A}~{}\mbox{wins}]\leq\Pr[\mathcal{B}~{}\mbox{wins}]+\eta^{2}.$ (35)
Because $f$ is a OWF, we have
$\Pr[\mathcal{A}~{}\mbox{wins}]\leq{\mathsf{negl}}(\lambda)+\eta^{2},$ (36)
which shows that the OWSG is $(\eta^{2}+{\mathsf{negl}}(\lambda))$-secure. From Lemma 2.13, the output length of the OWSG is at most $O(\log(\ell)\log(\eta^{-1}))$.
Now we show Equation 35. Let $\mathcal{A}$ be a QPT adversary of the OWSG. From this $\mathcal{A}$, we construct an adversary $\mathcal{B}$ of the OWF $f$ as follows:
  1. 1.
    Given $(1^{\lambda},f(k))$, run $k^{\prime}\leftarrow\mathcal{A}(1^{\lambda},\ket{h_{f(k)}}^{\otimes t})$.
  2. 2.
    Output $k^{\prime}$.
From Lemma 2.13, $|\langle h_{f(k)}|h_{f(k^{\prime})}\rangle|\leq\eta$ if $f(k)\neq f(k^{\prime})$. Therefore we have
$\Pr[\mathcal{A}~{}\mbox{wins}]$ $=\frac{1}{2^{\lambda}}\sum_{k}\sum_{k^{\prime}}\Pr[k^{\prime}\leftarrow\mathcal{A}(1^{\lambda},\ket{h_{f(k)}}^{\otimes t})]|\langle h_{f(k)}|h_{f(k^{\prime})}\rangle|^{2}$ (37)
$=\frac{1}{2^{\lambda}}\sum_{k}\sum_{k^{\prime}:f(k^{\prime})=f(k)}\Pr[k^{\prime}\leftarrow\mathcal{A}(1^{\lambda},\ket{h_{f(k)}}^{\otimes t})]|\langle h_{f(k)}|h_{f(k^{\prime})}\rangle|^{2}$ (38)
$+\frac{1}{2^{\lambda}}\sum_{k}\sum_{k^{\prime}:f(k^{\prime})\neq f(k)}\Pr[k^{\prime}\leftarrow\mathcal{A}(1^{\lambda},\ket{h_{f(k)}}^{\otimes t})]|\langle h_{f(k)}|h_{f(k^{\prime})}\rangle|^{2}$ (39)
$\leq\frac{1}{2^{\lambda}}\sum_{k}\sum_{k^{\prime}:f(k^{\prime})=f(k)}\Pr[k^{\prime}\leftarrow\mathcal{A}(1^{\lambda},\ket{h_{f(k)}}^{\otimes t})]$ (40)
$+\frac{1}{2^{\lambda}}\sum_{k}\sum_{k^{\prime}:f(k^{\prime})\neq f(k)}\Pr[k^{\prime}\leftarrow\mathcal{A}(1^{\lambda},\ket{h_{f(k)}}^{\otimes t})]\eta^{2}$ (41)
$\leq\frac{1}{2^{\lambda}}\sum_{k}\sum_{k^{\prime}:f(k^{\prime})=f(k)}\Pr[k^{\prime}\leftarrow\mathcal{A}(1^{\lambda},\ket{h_{f(k)}}^{\otimes t})]+\eta^{2}$ (42)
$\leq\Pr[\mathcal{B}~{}\mbox{wins}]+\eta^{2}.$ (43)
Theorem 1.6 follows from Theorem 4.5 as an immediate corollary.
Corollary 4.6 (Restatement of Theorem 1.6).
For any constant $\epsilon>0$, there exist $\epsilon$-OWSGs with $O(\log\log\lambda)$-qubit pure state outputs assuming the existence of subexponentially secure OWFs.
Proof.
Without loss of generality, we can assume that a subexponentially secure OWF is length-preserving. i.e., it maps $\lambda$-bit inputs to $\lambda$-bit outputs. Let $f$ be a length-preserving subexponentially secure OWFs. Since $f$ is subexponentially secure there is a constant $10<c<$ 1 such that for any QPT adversary $\mathcal{A}$,
$\Pr[f(x^{\prime})=f(x):x\leftarrow\{0,1\}^{\lambda},x^{\prime}\leftarrow\mathcal{A}(1^{\lambda},f(x))]\leq 2^{-\lambda^{c}}$ (44)
for all sufficiently large $\lambda$. Then we construct a OWF $g$ that maps $\lambda$-bit inputs to $\ell(\lambda):=\lfloor(\log\lambda)^{2c^{-1}}\rfloor$-bit outputs as follows: for any $x\in\{0,1\}^{\lambda}$, $g(x):=f(x_{[\ell]})$ where $x_{[\ell]}$ is the first $\ell$ bits of $x$. By a straightforward reduction, for any QPT adversary $\mathcal{A}$, we have
$\Pr[g(x^{\prime})=g(x):x\leftarrow\{0,1\}^{\lambda},x^{\prime}\leftarrow\mathcal{A}(1^{\lambda},g(x))]\leq 2^{-\ell^{c}}\leq 2^{-(\log\lambda)^{2}}={\mathsf{negl}}(\lambda).$ (45)
Thus, $g$ is a OWF that maps $\lambda$-bit inputs to $\ell(\lambda)=\lfloor(\log\lambda)^{2c^{-1}}\rfloor$-bit outputs. By applying Theorem 4.5 for $\eta:=(\epsilon/2)^{1/2}$, there is a $(\epsilon/2+{\mathsf{negl}}(\lambda))$-OWSG with $O(\log\ell)=O(\log\log\lambda)$-qubit pure state outputs. Since $\epsilon/2+{\mathsf{negl}}(\lambda)<\epsilon$ for sufficiently large $\lambda$, this implies Corollary 4.6. ∎
4.3 Upper Bound for Weak OWSGs
For proving Theorem 1.7, we prove the following theorem.
Theorem 4.7.
If there exist OWFs that map $\lambda$-bit inputs to $\ell(\lambda)$-bit outputs, then there exist weak OWSGs with $\lceil\frac{\ell(\lambda)}{\log\lambda}\rceil$-qubit outputs.
Proof.
We simply write $\ell$ to mean $\ell(\lambda)$ for notational simplicity. Let $f$ be a OWF that maps $\lambda$-bit inputs to $\ell$-bit outputs. Define $t:=\lceil\frac{\ell}{\log\lambda}\rceil$, and we embed $\{0,1\}^{\ell}$ into $[\lambda]^{t}$. Then we can regard $f$ as a function that maps $\lambda$-bit strings to an element of $[\lambda]^{t}$. For $\theta$, let
|$\ket{+_{\theta}}:=\frac{1}{\sqrt{2}}(\ket{0}+e^{i\theta}\ket{1}).$
We construct a OWSG as follows:
  • $\mathsf{KeyGen}(1^{\lambda})\to k:$ Sample $k\leftarrow\{0,1\}^{\lambda}$.
  • $\mathsf{StateGen}(k)\to\ket{\phi_{k}}:$ Compute $f(k)=(y_{1},y_{2},...,y_{t})\in[\lambda]^{t}$ and output |$\ket{\phi_{k}}\coloneqq\bigotimes_{j\in[t]}\ket{+_{2\pi y_{j}/\lambda}}$.
  • $\mathsf{Ver}(k^{\prime},\ket{\phi_{k}})\to\top/\bot:$ Project |$\ket{\phi_{k}}$ onto |$\ket{\phi_{k^{\prime}}}$. If the projection is successful, output $\top$. Otherwise, output $\bot$.
For any $y\neq y^{\prime}$, we have $|\langle+_{2\pi y/\lambda}|+_{2\pi y^{\prime}/\lambda}\rangle|\leq\sqrt{\frac{1+\cos(2\pi/\lambda)}{2}}$. This is because
$|\langle+_{2\pi y/\lambda}|+_{2\pi y^{\prime}/\lambda}\rangle|$ $=\left|\frac{1+e^{i2\pi(y-y^{\prime})/\lambda}}{2}\right|$ (46)
$=\left|\frac{1+\cos(2\pi(y-y^{\prime})/\lambda)+i\sin(2\pi(y-y^{\prime}))}{2}\right|$ (47)
$=\sqrt{\left(\frac{1+\cos(2\pi(y-y^{\prime})/\lambda)}{2}\right)^{2}+\frac{\sin^{2}(2\pi(y-y^{\prime}))}{4}}$ (48)
$\leq\sqrt{\frac{1+\cos(2\pi(y-y^{\prime})/\lambda)}{2}}$ (49)
$\leq\sqrt{\frac{1+\cos(2\pi/\lambda)}{2}}.$ (50)
Thus, for any $k,k^{\prime}$ such that $f(k)\neq f(k^{\prime})$, we have $|\langle\phi_{k}|\phi_{k^{\prime}}\rangle|\leq\sqrt{\frac{1+\cos(2\pi/\lambda)}{2}}$. By a similar reduction to that in the proof of Theorem 4.5, we can see that the above OWSG satisfies $\left(\frac{1+\cos(2\pi/\lambda)}{2}+{\mathsf{negl}}(\lambda)\right)$-one-wayness. Since $\cos\theta=1-\theta^{2}/2+O(\theta^{4})$, we have $\cos(2\pi/\lambda)\leq 1-(\pi/\lambda)^{2}$ for sufficiently large $\lambda$. Thus, $\frac{1+\cos(2\pi/\lambda)}{2}=1-\frac{1}{{\mathrm{poly}}(\lambda)}$ and thus the above is a weak OWSG. ∎
Theorem 1.7 follows from Theorem 4.7 as an immediate corollary.
Corollary 4.8 (Restatement of Theorem 1.7).
For any QPT-computable function $m(\lambda)=\omega(1)$, there exist weak OWSGs with $m(\lambda)$-qubit pure state outputs assuming the existence of exponentially secure OWFs that map $\lambda$-bit inputs to $O(\lambda)$-bit outputs.
Proof.
Suppose that $f$ is an exponentially secure OWF that maps $\lambda$-bit inputs to $k\lambda$-bit outputs for some constant $k\geq 1$. Since it is exponentially secure, there is a constant $10<c<$ 1 such that for any QPT adversary $\mathcal{A}$,
$\Pr[f(x^{\prime})=f(x):x\leftarrow\{0,1\}^{\lambda},x^{\prime}\leftarrow\mathcal{A}(1^{\lambda},f(x))]\leq 2^{-c\lambda}$ (51)
for all sufficiently large $\lambda$. Then we construct a OWF $g$ that maps $\lambda$-bit inputs to $\ell(\lambda):=\lfloor m(\lambda)\log\lambda\rfloor$-bit outputs as follows: for any $x\in\{0,1\}^{\lambda}$, $g(x):=f(x_{[\ell]})$ where $x_{[\ell]}$ is the first $\ell$ bits of $x$. By a straightforward reduction, for any QPT adversary $\mathcal{A}$, we have
$\Pr[g(x^{\prime})=g(x):x\leftarrow\{0,1\}^{\lambda},x^{\prime}\leftarrow\mathcal{A}(1^{\lambda},g(x))]\leq 2^{-c\ell}\leq 2^{c}\cdot\lambda^{-cm(\lambda)}={\mathsf{negl}}(\lambda),$ (52)
where the final equality follows from $m(\lambda)=\omega(1)$. Thus, $g$ is a OWF that maps $\lambda$-bit inputs to $\ell(\lambda)=\lfloor m(\lambda)\log\lambda\rfloor$-bit outputs. By applying Theorem 4.7, there is a weak OWSG with $\lceil\frac{\ell(\lambda)}{\log\lambda}\rceil$-qubit outputs. By the definition of $\ell(\lambda)$, we have $\lceil\frac{\ell(\lambda)}{\log\lambda}\rceil\leq m(\lambda)$. Since we can arbitrarily increase the output length of weak OWSGs by a trivial padding, this implies Corollary 4.8. ∎
5 EFIs
In this section, we show upper and lower bounds for the output length of EFIs.
5.1 Lower Bound
Theorem 5.1 (Restatement of Theorem 1.8).
There do not exist $O(\log\lambda)$-qubit EFIs.
Proof.
Let $\mathsf{StateGen}$ be a QPT algorithm that takes a bit $b$ and $1^{\lambda}$ for $\lambda\in\mathbb{N}$ as input and outputs an $\ell(\lambda)$-qubit state $\rho_{b}$ for $\ell(\lambda)=O(\log\lambda)$. Suppose that $\rho_{0}$ and $\rho_{1}$ are statistically distinguishable, i.e., $\|\rho_{0}-\rho_{1}\|_{tr}\geq\frac{1}{p(\lambda)}$ for some polynomial $p$. It suffices to show that $\rho_{0}$ and $\rho_{1}$ are not computationally indistinguishable. To show this, we consider the following adversary $\mathcal{A}$ that distinguishes $\rho_{0}$ and $\rho_{1}$.
  1. 1.
    Take the security parameter $1^{\lambda}$ and a quantum state $\sigma$, which is either of $\rho_{0}$ or $\rho_{1}$ as input.
  2. 2.
    Set $\delta=\delta(\lambda):=\frac{1}{16p(\lambda)}$.
  3. 3.
    For $b\in\{0,1\}$, generate $t:=144\lambda 2^{4\ell}/\delta^{2}$ copies of $\rho_{b}$ and run $M_{b}\leftarrow\mathsf{Tomography}(\rho^{\otimes t}_{b},\delta)$ where $\mathsf{Tomography}$ is the algorithm in Lemma 2.10.
  4. 4.
    Compute the spectral decomposition of $M_{0}-M_{1}$:
    $M_{0}-M_{1}=\sum_{i}\mu_{i}\ket{v_{i}}\bra{v_{i}}$ |
    where |$\ket{v_{i}}$ is an eigenvector with eigenvalue $\mu_{i}$ and {|$\{\ket{v_{i}}\}_{i}$ forms an orthonormal basis of $\mathbb{C}^{2^{\ell}}$. Define
    $Q:=\sum_{i:\mu_{i}\geq 0}\mu_{i}\ket{v_{i}}\bra{v_{i}},~{}~{}~{}S:=\sum_{i:\mu_{i}<0}-\mu_{i}\ket{v_{i}}\bra{v_{i}}$ |
    so that $Q$ and $S$ are positive semi-definite matrices with orthogonal supports and satisfies $M_{0}-M_{1}=Q-S$. Let $\Pi$ be the projector onto the support of $Q$, i.e.,
    $\Pi:=\sum_{i:\mu_{i}\geq 0}\ket{v_{i}}\bra{v_{i}}.$ | .
  5. 5.
    Apply projective measurement $\{\Pi,I-\Pi\}$ to $\sigma$ and output the measurement outcome.
We remark that $\mathcal{A}$ may not be QPT since there may not exist a polynomial size (or even finite size) quantum circuit that exactly implements the projective measurement $\{\Pi,I-\Pi\}$. However, we observe that we can approximate $\mathcal{A}$ in QPT with an exponentially small error as follows. First, tomography in Step 3 runs in QPT since $t=144\lambda 2^{4\ell}/\delta^{2}={\mathrm{poly}}(\lambda)$ and the running time of $\mathsf{Tomography}$ is polynomial in $1/\delta,d=2^{\ell}$ and $\lambda$ as required in Lemma 2.10. In Step 4, by a numerical implementation of a standard diagonalization algorithm, we can obtain a description of a projector $\widetilde{\Pi}$ (which consists of all the entries of the matrix $\widetilde{\Pi}$) such that $\|\Pi-\widetilde{\Pi}\|_{op}\leq 2^{-\Omega(\lambda)}$ in QPT where $\|\cdot\|_{op}$ means the operator norm. It is known that any $\ell$-qubit unitary can be implemented with an error $\epsilon$ (in terms of the operator norm) by a ${\mathrm{poly}}(2^{\ell},\log\epsilon^{-1})$-quantum circuit [NC10, Sec. 4.5]. By inspecting the proof, we can see that we can compute the description of the approximating quantum circuit in (classical) time ${\mathrm{poly}}(2^{\ell},\log\epsilon^{-1})$ given the description of the unitary. By setting $\epsilon:=2^{-\Omega(\lambda)}$, the above implies that we can compute a description of a quantum circuit that implements projective measurement $\{\widetilde{\Pi}^{\prime},I-\widetilde{\Pi}^{\prime}\}$ such that $\|\widetilde{\Pi}-\widetilde{\Pi}^{\prime}\|_{op}\leq 2^{-\Omega(\lambda)}$. By the triangle inequality, $\|\Pi-\widetilde{\Pi}^{\prime}\|_{op}\leq 2^{-\Omega(\lambda)}$, and thus the output distribution changes only by $2^{-\Omega(\lambda)}$ even if we apply $\{\widetilde{\Pi}^{\prime},I-\widetilde{\Pi}^{\prime}\}$ instead of $\{\Pi,I-\Pi\}$. This means that $\mathcal{A}$ can be approximated with an exponentially small error in QPT.
Thus, it suffices to show that $\mathcal{A}$’s distinguishing advantage is $1/{\mathrm{poly}}(\lambda)$. Moreover, since $\|M_{b}-\rho_{b}\|_{tr}\leq\delta$ holds for $b\in\{0,1\}$ except for a negligible probability by Lemma 2.10, it suffices to show that $\mathcal{A}$’s distinguishing advantage is $1/{\mathrm{poly}}(\lambda)$ conditioned on that $\|M_{b}-\rho_{b}\|_{tr}\leq\delta$ for $b\in\{0,1\}$. We show this below.
First, by the definitions of $Q$ and $S$, we have
$\|M_{0}-M_{1}\|_{tr}=\frac{\mathrm{Tr}(Q)+\mathrm{Tr}(S)}{2}.$ (53)
We also have
$\left|\frac{\mathrm{Tr}(Q)-\mathrm{Tr}(S)}{2}\right|$ $=\left|\frac{\mathrm{Tr}(M_{0})-\mathrm{Tr}(M_{1})}{2}\right|$ (54)
$\leq\frac{1}{2}\left(|\mathrm{Tr}(\rho_{0}-\rho_{1})|+|\mathrm{Tr}(M_{0}-\rho_{0})|+|\mathrm{Tr}(M_{1}-\rho_{1})|\right)$ (55)
$\leq 0+\|M_{0}-\rho_{0}\|_{tr}+\|M_{1}-\rho_{1}\|_{tr}$ (56)
$\leq 2\delta$ (57)
where Equation 54 follows from $Q-S=M_{0}-M_{1}$, Equation 55 follows from the triangle inequality, Equation 56 follows from $\mathrm{Tr}(\rho_{0})=\mathrm{Tr}(\rho_{1})=1$ and $\mathrm{Tr}(M)\leq\|M\|_{1}=2\|M\|_{tr}$ for any Hermitian matrix $M$, and Equation 57 follows from the assumption that $\|M_{b}-\rho_{b}\|_{tr}\leq\delta$ for $b\in\{0,1\}$. Then we have
$\mathrm{Tr}(\Pi(M_{0}-M_{1}))$ $=\mathrm{Tr}(Q)$ (58)
$=\frac{\mathrm{Tr}(Q)+\mathrm{Tr}(S)}{2}+\frac{\mathrm{Tr}(Q)-\mathrm{Tr}(S)}{2}$ (59)
$\geq\|M_{0}-M_{1}\|_{tr}-2\delta$ (60)
$\geq\|\rho_{0}-\rho_{1}\|_{tr}-\|M_{0}-\rho_{0}\|_{tr}-\|M_{1}-\rho_{1}\|_{tr}-2\delta$ (61)
$\geq\|\rho_{0}-\rho_{1}\|_{tr}-4\delta$ (62)
where Equation 58 follows from the definitions of $Q,S$, and $\Pi$, Equation 60 follows from Equation 53 and Equations 54, 55, 56 and 57, Equation 61 follows from the triangle inequality, and Equation 62 follows from the assumption that $\|M_{b}-\rho_{b}\|_{tr}\leq\delta$ for $b\in\{0,1\}$.
Then we have
$\mathrm{Tr}(\Pi\rho_{0})-\mathrm{Tr}(\Pi\rho_{1})$ $=\mathrm{Tr}(\Pi(M_{0}-M_{1}))+\mathrm{Tr}(\Pi(\rho_{0}-M_{0}))+\mathrm{Tr}(\Pi(M_{1}-\rho_{1}))$ (63)
$\geq\|\rho_{0}-\rho_{1}\|_{tr}-4\delta-2\|M_{0}-\rho_{0}\|_{tr}-2\|M_{1}-\rho_{1}\|_{tr}$ (64)
$\geq\|\rho_{0}-\rho_{1}\|_{tr}-8\delta$ (65)
$=\frac{1}{2p(\lambda)}$ (66)
where Equation 64 follows from Equations 58, 59, 60, 61 and 62 and Corollary 2.16, Equation 65 follows from the assumption that $\|M_{b}-\rho_{b}\|_{tr}\leq\delta$ for $b\in\{0,1\}$, and Equation 66 follows from $\|\rho_{0}-\rho_{1}\|_{tr}\geq\frac{1}{p(\lambda)}$ and $\delta=\frac{1}{16p(\lambda)}$. ∎
5.2 Upper Bound
We show a matching upper bound assuming the existence of exponentially secure PRGs.
Theorem 5.2 (Restatement of Theorem 1.9).
If exponentially secure PRGs exist, then $m(\lambda)$-qubit EFI pairs exist for any QPT computable function $m(\lambda)=\omega(\log\lambda)$.
Proof.
Let $G$ be an exponentially secure PRG. Without loss of generality, we assume that $G$ is length-doubling, i.e., it maps $n$-bit strings to $2n$-bit strings. [^10] This is because we can stretch the output length by sequential application which preserves the exponential security. The exponential security means that there exists a constant $10<c<$ 1 such that for any QPT adversary $\mathcal{A}$
$\left|\Pr_{x\leftarrow\{0,1\}^{\lambda}}[1\leftarrow\mathcal{A}(1^{\lambda},G(x))]-\Pr_{y\leftarrow\{0,1\}^{2\lambda}}[1\leftarrow\mathcal{A}(1^{\lambda},y)]\right|\leq 2^{-c\lambda}$ (67)
for all sufficiently large $\lambda$. Let $n(\lambda):=\lceil m(\lambda)/2\rceil$. Then we have
$\left|\Pr_{x\leftarrow\{0,1\}^{n}}[1\leftarrow\mathcal{A}(1^{n},G(x))]-\Pr_{y\leftarrow\{0,1\}^{2n}}[1\leftarrow\mathcal{A}(1^{n},y)]\right|\leq 2^{-cn}={\mathsf{negl}}(\lambda)$ (68)
for sufficiently large $\lambda$. Define EFI pairs as follows:
$\rho_{0}$ $\coloneqq\frac{1}{2^{n}}\sum_{x\in\{0,1\}^{n}}\ket{G(x)}\bra{G(x)},$ (69)
$\rho_{0}$ $\coloneqq\frac{1}{2^{2n}}\sum_{y\in\{0,1\}^{2n}}\ket{y}\bra{y}$ | (70)
The computational indistinguishability immediately follows from Equation 68. On the other hand, Lemma 2.17 implies
$F(\rho_{0},\rho_{1})\leq 2^{-n}={\mathsf{negl}}(\lambda)$ (71)
and thus $\|\rho_{0}-\rho_{1}\|_{tr}=1-{\mathsf{negl}}(\lambda)$. Thus, they are stastically far. Finally, $\rho_{0}$ and $\rho_{1}$ are $2n$ qubit states. When $m$ is even, then $2n=m$ and thus it gives $m$-qubit EFI pairs. When $m$ is odd, then $12n=m-$ 1, but we can add append one-qubit state |$\ket{0}$ to both $\rho_{0}$ and $\rho_{1}$ so that they are $m$-qubit states while preserving computational indistinguishability and statistical farness. ∎
Acknowledgements. This work was initiated while the authors were joining NII Shonan Meeting No.198 “New Directions in Provable Quantum Advantages”. We thank Ryu Hayakawa and Fermi Ma for useful conversations. We thank anonymous reviewers of CRYPTO 2024 and TQC 2024 for their valuable comments, especially for pointing out that injectivity is not needed in Theorem 4.7. MH is supported by a KIAS Individual Grant QP089801. TM is supported by JST CREST JPMJCR23I3, JST Moonshot R&D JPMJMS2061-5-1-1, JST FOREST, MEXT QLEAP, the Grant-in Aid for Transformative Research Areas (A) 21H05183, and the Grant-in-Aid for Scientific Research (A) No.22H00522.
Appendix A Quantum Bit Commitments
A.1 Preliminaries
Definition A.1 (Canonical Quantum Bit Commitments [Yan22]).
A canonical quantum bit commitment scheme is a family $\{Q_{0}(\lambda),Q_{1}(\lambda)\}_{\lambda\in{\mathbb{N}}}$ of efficiently implementable unitary operators that is used as follows.
  • Commit phase:
    In the commit phase, the sender who wants to commit $b\in\{0,1\}$ generates the state |$\ket{\Psi_{b}}_{\mathbf{R},\mathbf{C}}\coloneqq Q_{b}(\lambda)|0...0\rangle$ over two registers $\mathbf{R}$ and $\mathbf{C}$. The sender sends the commitment register $\mathbf{C}$ to the receiver.
  • Reveal phase:
    In the reveal phase, the sender sends the reveal register $\mathbf{R}$ and $b$ to the receiver. The receiver applies $Q_{b}^{\dagger}(\lambda)$ on the state over $\mathbf{R}$ and $\mathbf{C}$, and measures all qubits in the computational basis. If all results are 0, the receiver accepts $b$. Otherwise, the receiver rejects.
For the notational simplicity, we often omit the security parameter $\lambda$ of $\{Q_{0}(\lambda),Q_{1}(\lambda)\}_{\lambda\in{\mathbb{N}}}$, and just write as $(Q_{0},Q_{1})$.
Definition A.2 (Hiding).
We say that a commitment scheme $(Q_{0},Q_{1})$ is statistically (resp. computationally) hiding if
$|\Pr[1\leftarrow\mathcal{A}(1^{\lambda},\rho_{0})]-\Pr[1\leftarrow\mathcal{A}(1^{\lambda},\rho_{1})]|\leq{\mathsf{negl}}(\lambda)$ (72)
for any unbounded (resp. QPT) adversary $\mathcal{A}$, where $\rho_{b}\coloneqq\mathrm{Tr}_{\mathbf{R}}(\ket{\Psi_{b}}_{\mathbf{R},\mathbf{C}})$, and $\mathrm{Tr}_{\mathbf{A}}(\sigma_{\mathbf{A},\mathbf{B}})$ means tracing out the register $\mathbf{A}$ of the state $\sigma_{\mathbf{A},\mathbf{B}}$ over two registers $\mathbf{A}$ and $\mathbf{B}$.
Definition A.3 (Binding).
We say that a commitment scheme $(Q_{0},Q_{1})$ is statistically (resp. computationally) binding if the winning probability of the following security game is ${\mathsf{negl}}(\lambda)$ for any unbounded (resp. QPT) adversary $\mathcal{A}$:
  1. 1.
    $\mathcal{A}$ (who plays the role of the sender) honestly commits 0.
  2. 2.
    $\mathcal{A}$ applies any operation on the register $\mathbf{R}$.
  3. 3.
    $\mathcal{A}$ opens the commitment to 1.
  4. 4.
    $\mathcal{A}$ wins if and only if the receiver accepts 1.
Remark A.4.
The above definition of binding, so-called honest binding, does not capture all possible attacks, because we assume that the adversary honestly commits 0 in the commit phase. However, when we want to show lower bounds, restricting to the honest binding only strengthens our results, because any reasonable definition of binding should satisfy the honest binding. Moreover, it is shown in [Yan22] that the honest binding implies a standard definition of binding, so-called sum-binding [DMS00], which captures general attacks. In the following, if we just say binding, it means honest binding.
Remark A.5.
In the original definition of canonical bit commitments in [Yan22], computational hiding and computational binding are required to hold against non-uniform QPT adversaries whereas we only require them against uniform QPT adversaries. This only makes our result on the lower bound for the commitment length stronger.
Lemma A.6 (Flavor Conversion [HMY23]).
Let $(Q_{0},Q_{1})$ be a statistically hiding (resp. binding) and computationally binding (resp. hiding) non-interactive quantum bit commitment scheme. Then, $(\tilde{Q}_{0},\tilde{Q}_{1})$ defined as follows is a statistically-binding (resp. hiding) and computationally-hiding (resp. binding) non-interactive quantum bit commitment scheme.
|$\ket{\tilde{\Psi}_{b}}_{\mathbf{R}^{\prime},\mathbf{C}^{\prime}}\coloneqq\tilde{Q}_{b}|0...0\rangle\coloneqq\frac{1}{\sqrt{2}}\left[|0\rangle_{\mathbf{D}}\otimes(Q_{0}|0...0\rangle)_{\mathbf{R},\mathbf{C}}+(-1)^{b}|1\rangle_{\mathbf{D}}\otimes(Q_{1}|0...0\rangle)_{\mathbf{R},\mathbf{C}}\right]$ (73)
for each $b\in\{0,1\}$, where $\mathbf{R}^{\prime}\coloneqq\mathbf{C}$ and $\mathbf{C}^{\prime}\coloneqq(\mathbf{R},\mathbf{D})$.
Lemma A.7.
For any two states $\rho$ and $\sigma$, $\mathrm{Tr}(\rho\sigma)\leq F(\rho,\sigma)$, where $F$ is the fidelity.
Proof.
The operator $\sqrt{\sigma}\rho\sqrt{\sigma}$ is positive semi-definite. Therefore, it has the diagonalization $\sqrt{\sigma}\rho\sqrt{\sigma}=\sum_{i}\alpha_{i}|\phi_{i}\rangle\langle\phi_{i}|$, where $\alpha_{i}\geq 0$ for all $i$. Hence
$F(\rho,\sigma)=\left(\mathrm{Tr}\sqrt{\sqrt{\sigma}\rho\sqrt{\sigma}}\right)^{2}=\left(\sum_{i}\sqrt{\alpha_{i}}\right)^{2}=\sum_{i}\alpha_{i}+\sum_{i\neq j}\sqrt{\alpha_{i}\alpha_{j}}\geq\sum_{i}\alpha_{i}=\mathrm{Tr}(\rho\sigma).$ (74)
A.2 Lower Bound
Theorem A.8.
There do not exist computationally-binding and computationally-hiding canonical quantum bit commitments with $O(\log\lambda)$-qubit commitment register.
By the flavor conversion (Lemma A.6), the above theorem implies the following corollary.
Corollary A.9.
There do not exist computationally-binding and computationally-hiding canonical quantum bit commitments with $O(\log\lambda)$-qubit reveal register.
Below, we prove Theorem A.8.
Proof of Theorem A.8.
[…]
  1. 1.
    There exists a polynomial $p$ such that $\left\|\rho_{0}-\rho_{1}\right\|_{tr}\geq\frac{1}{p(\lambda)}$ for infinitely many $\lambda$.
  2. 2.
    $\left\|\rho_{0}-\rho_{1}\right\|_{tr}\leq{\mathsf{negl}}(\lambda)$.
In the first case, we can show that the scheme does not satisfy computational hiding by exactly the same proof as that of Theorem 5.1. The only difference is that our assumption is $\left\|\rho_{0}-\rho_{1}\right\|_{tr}\geq\frac{1}{p(\lambda)}$ for infinitely many $\lambda$ instead of for all $\lambda$. Thus, by repeating exactly the same argument as in the proof of Theorem 5.1, we can obtain a QPT adversary that distinguishes $\rho_{0}$ and $\rho_{1}$ with advantage at least $\frac{1}{2p(\lambda)}$ for infinitely many $\lambda$. This means that the scheme does not satisfy computational hiding.
In the second case, the scheme satisfies statistical hiding. We prove that it cannot satisfy computational binding. Toward contradiction, we assume that it satisfies computational binding. Then, due to the flavor conversion (Lemma A.6), we have a statistically-binding and computationally-hiding canonical quantum bit commitment scheme $\{\tilde{Q}_{0}(\lambda),\tilde{Q}_{1}(\lambda)\}_{\lambda\in\mathbb{N}}$ with $|\mathbf{R}^{\prime}|=|\mathbf{C}|=O(\log\lambda)$, where $|\mathbf{A}|$ is the number of qubits of the register $\mathbf{A}$. However, we can show that such commitments do not exist, because we can construct a QPT adversary that breaks the computational hiding as follows:
  1. 1.
    Given $\tilde{\rho}_{b}\coloneqq\mathrm{Tr}_{\mathbf{R}^{\prime}}(\ket{\tilde{\Psi}_{b}}_{\mathbf{R}^{\prime},\mathbf{C}^{\prime}})$ as input.
  2. 2.
    Run the swap test between $\tilde{\rho}_{0}$ and $\tilde{\rho}_{b}$.
  3. 3.
    Output the output of the swap test.
Then the probability $p_{b}$ that the adversary outputs 1 is $p_{0}=\frac{1+\mathrm{Tr}(\tilde{\rho}_{0}^{2})}{2}$ and $p_{1}=\frac{1+\mathrm{Tr}(\tilde{\rho}_{0}\tilde{\rho}_{1})}{2}$. Because of the statistical binding, we have $F(\tilde{\rho}_{0},\tilde{\rho}_{1})\leq{\mathsf{negl}}(\lambda)$. Therefore, by using Lemma A.7,
$\mathrm{Tr}(\tilde{\rho}_{0}\tilde{\rho}_{1})\leq F(\tilde{\rho}_{0},\tilde{\rho}_{1})\leq{\mathsf{negl}}(\lambda).$ (75)
On the other hand, let $\tilde{\rho}_{0}=\sum_{i\in[N]}p_{i}\ket{v_{i}}\bra{v_{i}}$ | be the spectral decomposition of $\tilde{\rho}_{0}$ where $N\leq 2^{|\mathbf{R}^{\prime}|}={\mathrm{poly}}(\lambda)$. Then
$\mathrm{Tr}(\tilde{\rho}_{0}^{2})=\sum_{i\in[N]}p_{i}^{2}\geq\frac{\left(\sum_{i\in[N]}p_{i}\right)^{2}}{N}=\frac{1}{{\mathrm{poly}}(\lambda)}$ (76)
where we used the Cauchy-Schwarz inequality. Therefore the QPT adversary can break the computational hiding. Thus, the commitment scheme $\{Q_{0}(\lambda),Q_{1}(\lambda)\}_{\lambda\in\mathbb{N}}$ does not satisfy computational binding.
In summary, $\{Q_{0}(\lambda),Q_{1}(\lambda)\}_{\lambda\in\mathbb{N}}$ does not satisfy computational hiding in the first case and does not satisfy computational binding in the second case. This completes the proof of Theorem A.8. ∎
Remark A.10.
It is known that any (even interactive) quantum commitment schemes can be converted into canonical one [Yan22], but the transformation could increase the commitment size. (Honest receiver’s register in the original commitment scheme is contained in the commitment register of the resulting one.) Therefore Theorem A.8 does not directly exclude interactive quantum commitments with $O(\log\lambda)$-qubit communication. It is an open problem whether interactive quantum commitments with $O(\log\lambda)$-qubit communication are possible or not.
A.3 Upper Bound
Similarly to the case of EFI pairs, we show a matching upper bound assuming the existence of exponentially secure PRGs.
Theorem A.11.
If exponentially secure PRGs exist, then statistically (resp. computationally) binding and computationally (resp. statistically) hiding canonical quantum bit commitments with $m(\lambda)$-qubit commitment register exist for any QPT computable function $m(\lambda)=\omega(\log\lambda)$.
Proof.
The proof for the case of statistically binding and computationally hiding commitments is similar to that of Theorem 5.2. Let $G$ be an exponentially secure PRG. Without loss of generality, we assume that $G$ is length-doubling. Let $n(\lambda):=\lceil m(\lambda)/2\rceil$ and define the canonical quantum bit commitment $(Q_{0},Q_{1})$ as follows:
|$\ket{\Psi_{0}}_{\mathbf{R},\mathbf{C}}$ $\coloneqq\frac{1}{\sqrt{2^{n}}}\sum_{x\in\{0,1\}^{n}}|x0^{n}\rangle_{\mathbf{R}}\otimes|G(x)\rangle_{\mathbf{C}},$ (77)
|$\ket{\Psi_{1}}_{\mathbf{R},\mathbf{C}}$ $\coloneqq\frac{1}{\sqrt{2^{2n}}}\sum_{y\in\{0,1\}^{2n}}|y\rangle_{\mathbf{R}}\otimes|y\rangle_{\mathbf{C}}.$ (78)
Its computational hiding follows from exponential security of $G$ and $n=\omega(\log\lambda)$. For its statistical binding, from Lemma 2.17, we have
$F(\rho_{0},\rho_{1})\leq 2^{-n}={\mathsf{negl}}(\lambda),$ (79)
which implies statistical binding (e.g., see [MY22b]). The number of qubits of the commitment register is $2n$. If $m$ is even, $2n=m$ and if $m$ is odd, $12n=m-$ 1, so we can simply append one-qubit to the commitment register to make it $m$-qubit.
The case of computationally binding and statistically hiding commitments immediately from the above and flavor conversion (Lemma A.6) noting that $\mathbf{C}$ and $\mathbf{R}$ have the same length in the above construction. ∎